### v2

# A Unified Study on Sequentiality in Universal Classification with Empirically Observed Statistics

Abstract
In the binary hypothesis testing problem, it is well known that sequentiality in taking samples eradicates the trade-off between two error exponents, yet implementing the optimal test requires the knowledge of the underlying distributions, say $P_{0}$ and $P_{1}$. In the scenario where the knowledge of distributions is replaced by empirically observed statistics from the respective distributions, the gain of sequentiality is less understood when subject to universality constraints over all possible $P_{0},P_{1}$. In this work, the gap is mended by a unified study on sequentiality in the universal binary classification problem, where the universality constraints are set on the expected stopping time as well as the type-I error exponent. The type-I error exponent is required to achieve a pre-set distribution-dependent constraint $\lambda(P_{0},P_{1})$ for all $P_{0},P_{1}$. Under the proposed framework, different sequential setups are investigated so that fair comparisons can be made with the fixed-length counterpart. By viewing these sequential classification problems as special cases of a general sequential composite hypothesis testing problem, the optimal type-II error exponents are characterized. Specifically, in the general sequential composite hypothesis testing problem subject to universality constraints, upper and lower bounds on the type-II error exponent are proved, and a sufficient condition for which the bounds coincide is given. The results for sequential classification problems are then obtained accordingly. With the characterization of the optimal error exponents, the benefit of sequentiality is shown both analytically and numerically by comparing the sequential and the fixed-length cases in representative examples of type-I exponent constraint $\lambda$.
I Introduction
It is known that sequentiality in taking samples greatly enhances reliability in statistical inference. Take the binary hypothesis testing problem as an example. The decision maker observes a sequence of samples drawn i.i.d. from one of the two known distributions $P_{0}$ or $P_{1}$. It aims to infer from which of the two distributions the sequence is generated. Both type-I and type-II error probabilities vanish exponentially fast as the number of samples $n$ tends to infinity. The exponential rates are denoted as the error exponents $e_{0}$ and $e_{1}$ respectively, between which there exists a fundamental trade-off [1]. When samples are taken sequentially and the decision maker is free to decide when to stop as long as the expected stopping time is less than a given constraint $n$, Wald’s Sequential Probability Ratio Test (SPRT) [2] is shown to be optimal [3] and the error exponents can simultaneously achieve the two extremes, namely, the two KL divergences $\mathrm{D}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$ and $\mathrm{D}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$. In words, sequentiality in taking samples eradicates the trade-off between error exponents. The optimal design of the stopping and inference strategies, however, critically relies on the knowledge of the underlying distributions [2, 3, 4].
In this work, we aim to investigate the benefit of sequentiality when the underlying distributions are unknown to the decision maker. Instead, the decision maker only has access to empirical statistics of training sequences sampled from these distributions, say, $P_{0}$ and $P_{1}$ in the binary-hypothesis case. Since the underlying distributions are unknown, it is natural to ask for a universal guarantee on certain performances. One such framework focused on the asymptotic performance as the number of samples in all sequences tend to infinity was proposed and studied by Ziv [5], where a universality constraint was set on the type-I error exponent to be no less than a given constant $\lambda_{0}>0$ regardless of the underlying distributions. Gutman [6] later proved the asymptotic optimality of Ziv’s universal test in the sense that it attains the optimal type-II error exponent (as a function of $P_{0},P_{1},\lambda_{0}$). Further extension was made by Levitan and Merhav [7] where a competitive criterion was taken, replacing the constant constraint $\lambda_{0}$ by a distribution-dependent one, $\lambda(P_{0},P_{1})$. From these results in the fixed-length settings, it can be seen that there also exists a trade-off between the two error exponents. A natural question emerges: in this universal classification problem, can sequentiality improve the trade-off, and even eradicate them without knowing the underlying distributions?
Full resolution to the question above is provided in this paper. Since the decision maker observes three sequences (one testing and two training), there are different ways of considering sequentiality. Main focus is on the following setups: (1) the $\mathsf{Fully\text{-}Sequential}$ setup where testing and training samples all arrive sequentially, (2) the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup where the training sequences have fixed length and testing samples arrive sequentially, and (3) the $\mathsf{Semi\text{-}Sequential\text{-}2}$ setup where the testing sequence has fixed length and training samples arrive sequentially. To have a fair comparison with the $\mathsf{Fixed\text{-}Length}$ setup in previous works, universality constraints are set on the expected stopping time and the type-I error exponent.
Our main contribution is the characterization of the optimal type-II error exponent when the type-I error exponent universality constraint $\lambda(P_{0},P_{1})$ satisfies mild continuity conditions. For the $\mathsf{Fully\text{-}Sequential}$ setup, the optimal type-II error exponent is $e_{1,\mathsf{seq}}^{*}(P_{0},P_{1})=\min\big{\{}\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\kappa(P_{0},P_{1})\big{\}}$. Here the first term $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$ is the Rényi divergence and $\alpha$ is the expected ratio of the number of the $P_{0}$-training samples to that of testing samples. It corresponds to the error event that the test stops at around time $n$ with testing and $P_{0}$-training samples looking like they are generated by the same distribution. Likewise, the second term $\kappa(P_{0},P_{1})$ corresponds to the error event that the test stops at around time $n$ while the testing samples look like $P_{1}$-training samples instead but the empirical distributions fall in the decision region of the null hypothesis. For the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup, there is an additional term as the performance is further restricted by the fixed-length training sequences. Specifically, the optimal type-II error exponent is $e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})=\min\big{\{}e_{1,\mathsf{seq}}^{*}(P_{0},P_{1}),\mu(P_{0},P_{1})\big{\}}$. The additional term $\mu(P_{0},P_{1})$ corresponds to an error event in the regime where the number of testing samples is much larger than the fixed lengths of training sequences and is hence the optimal type-II error exponent in an induced fixed-length composite hypothesis testing problem [7] when the distribution of the testing samples is fixed as $P_{1}$ and known. Similarly, for the $\mathsf{Semi\text{-}Sequential\text{-}2}$ setup, the additional restriction is from the fixed-length testing sequence, and $e_{1,\mathsf{semi2}}^{*}(P_{0},P_{1})=\min\big{\{}e_{1,\mathsf{seq}}^{*}(P_{0},P_{1}),\nu(P_{0},P_{1})\big{\}}$. Here $\nu(P_{0},P_{1})$ instead corresponds to an error event in the regime where the numbers of training samples are both much larger than the fixed length of the testing sequence so that effectively the distributions of the two hypotheses are both known. Hence, it is the optimal type-II error exponent for the binary hypothesis problem when the type-I error exponent is $\lambda(P_{0},P_{1})$. For the $\mathsf{Fixed\text{-}Length}$ setup, the optimal type-II error exponent is given in [7] and denoted as $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$. To illustrate the benefit of sequentiality and the optimal type-II error exponents in different setups, a numerical example is given in Figure 1, where $\mathsf{Fully\text{-}Sequential}$ strictly outperforms $\mathsf{Semi\text{-}Sequential\text{-}1}$/$\mathsf{Semi\text{-}Sequential\text{-}2}$, and $\mathsf{Semi\text{-}Sequential\text{-}1}$/$\mathsf{Semi\text{-}Sequential\text{-}2}$ strictly outperforms $\mathsf{Fixed\text{-}Length}$. Moreover, $\mathsf{Semi\text{-}Sequential\text{-}1}$ is sometimes better than $\mathsf{Semi\text{-}Sequential\text{-}2}$ and sometimes worse.
Figure 1: The optimal type-II error exponents under type-I error exponent constraint $\lambda(P_{0},P_{1})=\xi\big{(}\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)+0.003\big{)}$. Here $\mathcal{X}=\{0,1\}$, $\alpha=0.38$, $\beta=0.6$, and $P_{0}^{*}=[0.6,0.4]$, $P_{1}^{*}=[0.1,0.9]$. Let $\xi$ increase from $0.001$ to $1$ to obtain a curve. The optimal type-II error exponents in different setups can be identified by taking the minimum of the corresponding terms. This particular example is chosen judiciously so that all terms are active in the considered regime.
Analytical comparisons are also made besides numerical evaluation. In particular, for the choice of $\lambda(P_{0},P_{1})$ being a constant $\lambda_{0}>0$, we show that $\mathsf{Fixed\text{-}Length}$ tests have strictly smaller type-II error exponent than $\mathsf{Fully\text{-}Sequential}$ ones. Moreover, $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ tests have the same optimal type-II error exponents, indicating that there is no additional gain due to sequentiality in taking training samples when testing samples are sequentially observed. On the other hand, when the choice of $\lambda(P_{0},P_{1})$ permits exponentially vanishing error probabilities for all $P_{0},P_{1}$, it is shown that the trade-off between error exponents is eradicated in the $\mathsf{Fully\text{-}Sequential}$ case, demonstrating strict error exponent gain over $\mathsf{Fixed\text{-}Length}$ tests. Moreover, regarding the additional benefit of sequentiality in taking training samples when testing samples are already sequential, we characterize the necessary and sufficient condition of whether or not there exists a strict gap between the optimal exponents of $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ tests. The condition pertains to $\alpha$ and $\beta$, the expected ratios of the number of the $P_{0}/P_{1}$-training samples to that of the testing samples, respectively.
A unified proof of the optimal type-II error exponents in the classification problem is developed by investigating a more general composite hypothesis testing problem with $s$ independent sequences, each of which is either fixed-length or sequentially observed. In this problem, there are two disjoint composite hypothesis sets, say $\mathscr{P}_{0}$ and $\mathscr{P}_{1}$. Each of them consists of some $s$-tuples of distributions, where each element of an $s$-tuple is the distribution of one of the $s$ sequences. In the following, we may refer to “$s$-tuples of distributions” as just distributions for simplicity. In other words, $\mathscr{P}_{0}$ and $\mathscr{P}_{1}$ collect the possible underlying distributions of the sequences under two hypotheses. When all sequences are fixed-length, the problem was investigated in [7], and hence the proposed problem can be viewed as a generalization of that in [7] with sequential components. Similar to the setup in the classification problem, the universality constraints are set on the expected stopping time and the type-I error exponent. For this general composite hypothesis testing problem with sequential components, we establish upper and lower bounds on the type-II error exponents, as well as a sufficient condition such that the upper and lower bounds coincide. In particular, the sufficient condition pertains to whether or not the closure of the intersection of a particular closed set of $s$-tuples of distributions and $\mathscr{P}_{0}$ is the same as the intersection of that set and the closure of $\mathscr{P}_{0}$. That particular closed set collects all $s$-tuples of distributions with the distributions of the sequentially observed samples being fixed. The optimal type-II error exponents of the classification problem can be obtained by identifying the different sequential setups as special cases of the composite hypothesis testing problem. To be more specific, let $s=3$ and $\mathscr{P}_{0}$/$\mathscr{P}_{1}$ collect all the $3$-tuples of distributions such that the distribution of the testing sequence matches that of the $P_{0}$/$P_{1}$-training sequence and different from that of the $P_{1}$/$P_{0}$-training sequence. The three sequences are set to be fixed-length or sequentially observed according to the different setups. With the above specialization, we show that the aforementioned sufficient condition holds in these cases, and characterize the the optimal type-II error exponents in the classification problems.
In general, the type-II error exponent of the composite hypothesis testing problem can be viewed as the minimum of three terms, where each term is the KL divergence of a set of distributions from the real underlying distribution. For the converse part, the first term corresponds to the set $\mathscr{P}_{0}$, and the proof involves standard arguments based on data processing inequality and Wald’s identity, similar to [8]. The second term corresponds to the subset of distributions in $\mathscr{P}_{1}$ that are within a certain threshold from some distributions in $\mathscr{P}_{0}$. Specifically, the threshold is given by the type-I error exponent constraint function $\lambda$. The idea behind these two terms is that if the empirical distributions of the observed sequences are close to the real underlying distributions, then in order to satisfy the universality constraint on the expected stopping time, the test should not stop too late as this event happens with high probability. However, since the real underlying distributions are unknown, the test should stop early as long as the empirical distributions are close to any possible distributions in $\mathscr{P}_{0}$ or $\mathscr{P}_{1}$. Moreover, if the test stops at some time around time $n$, it should follow the optimal decision rule in the fixed-length setup [7] in order to satisfy the universality constraint on the type-I error exponent. That is, output 00 if the empirical distributions are within the threshold $\lambda$ from some distributions in $\mathscr{P}_{0}$. The main technique to make these arguments rigorous is the method of types. Lastly, the third term is the restriction caused by fixed-length sequences, if any, and is obtained by a reduction from a new fixed-length composite hypothesis testing problem where the distributions behind sequentially observed samples are known. Apparently the decision maker cannot do better than when it exactly knows those distributions.
For the achievability part, we propose a two-phase test that satisfies the universality constraints. Following the same idea as in the converse part, to ensure the expected stopping time is bounded by $n$, the test stops at time $n-1$ if the empirical distributions of the observed sequences are close to any distributions in $\mathscr{P}_{0}$ or $\mathscr{P}_{1}$. The decision rule follows the optimal fixed-length test in [7] to meet the type-I error exponent constraint. The contribution to the type-II error exponent in this part unsurprisingly corresponds to the first two terms. On the other hand, if the empirical distributions are not close to any possible underlying distributions, it means that the samples are inadequate to make a good estimation. Again, motivated by the idea in the converse part, we aim to approach the case when the distributions behind sequentially observed samples are known. Intuitively, if we observe infinitely many samples, then the distributions behind those sequences can be known by the law of large numbers. Thus in this case the test continues to take more samples until time $n^{2}$, which can be viewed as mimicking having infinitely many samples. The decision rule again leverages the idea of fixed-length test in [7]. In this part, the contribution to the type-II error exponent makes up the third term under some technical conditions to ensure convergence.
I-A Related works
For the sequential version of binary classification problem, Haghifam et al. [9] considered the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup as well. They proposed a test and showed that it achieves larger Bayesian error exponent over the $\mathsf{Fixed\text{-}Length}$ case. Under the same setting, Bai et al. [10] proposed an almost fixed-length two-phase test with performance lying between Gutman’s $\mathsf{Fixed\text{-}Length}$ test and the $\mathsf{Semi\text{-}Sequential\text{-}1}$ test in [9]. However, in both [9] and [10], they did not have a universality constraint on the expected stopping time, nor other universal guarantees over all possible distributions. For example, in [9], the expected stopping time of their test depends implicitly on the unknown distributions $P_{0},P_{1}$. Moreover, in order to achieve certain performance guarantees, parameters have to be chosen to satisfy some conditions that depend on the underlying distribution. A problem of lacking universality guarantees is that there may be a test that performs well under a specific pair of distributions and bad under another pair of distributions, while another test does the opposite. Then it cannot be said which test is better. Even if a test may be instance optimal, there might not exist a universally optimal test. As a result, it would be difficult or even impossible to derive a tight converse bound matching the achievability results. Such difficulties are experienced in both [9] and [10]. In comparison, by setting universality constraints, we can focus on tests that are (at least to certain extent) comparable to each other. For such tests, optimality can be clearly defined and there even exist tests that are universally optimal.
Hsu et al. [8] considered the same $\mathsf{Fully\text{-}Sequential}$ setup as in this work, while the universality constraints are set on the expected stopping time and vanishing error probabilities. Hence the result is not directly comparable with those works in the $\mathsf{Fixed\text{-}Length}$ setting [5, 6, 7]. In a preliminary version of this work [11], the treatment only pertains to universal tests that can achieve exponentially vanishing error probabilities, and thus faces a similar problem as [8], not comparable with results under a constant type-I error exponent constraint [5, 6]. Besides, the lengths of training sequences are assumed to be the same in [11], while in this work, the assumption is lifted. Partial results of this work also appear in [12], where the focus is on $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ classification. Here the results are generalized to a composite hypothesis testing problem and the proofs are unified. This provides the flexibility to consider more possibilities regarding the sequentiality in taking samples, and the potential to apply to other classical problems. For the sake of completeness, results in [11, 12] will also be included in exposition.
I-B Organization
In Section II, we formally define the binary classification problem, including different sequential setups, the universality constraints, and the performance metrics. Section III contains the main results for binary classification, namely, the optimal error exponents and comparison between different sequential setups. In Section IV we extend to a general composite hypothesis testing problem and show that the binary classification is a special case. Section V and Section VI contain respectively the proof of converse and achievability for the general results. In Section VII, we discuss some assumptions and straightforward extensions. Finally, Section VIII concludes the paper with some possible future directions.
I-C Notation
A finite-length sequence $(x_{1},x_{2},...,x_{n})$ is denoted as $x^{n}$. Logarithms are of base $2$ if not specified. For a subset $S$ of a topological space, $\mathrm{int}(S)$ and $\mathrm{cl}(S)$ denote the interior and closure of $S$, respectively. $\mathcal{P}(\mathcal{X})$ is the set of all probability distributions over alphabet $\mathcal{X}$. When $|\mathcal{X}|=d$, the probability simplex $\mathcal{P}(\mathcal{X})$ can be embedded into $\mathbb{R}^{d-1}$ and the Euclidean norm is used if not specified. Hence we use $\mathrm{int}(\mathcal{P}(\mathcal{X}))$ to denote the set of all probability distributions over alphabet $\mathcal{X}$ with full support. An indicator function is written as $\mathbbm{1}{\left\{\cdot\right\}}$. Given positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}\doteq b_{n}$ if $\lim_{n\to\infty}\frac{1}{n}\log\frac{a_{n}}{b_{n}}=0$. The relation $\mathrel{\dot{\leq}}$ is defined similarly. Also, $o(1)$ is used to denote vanishing terms, and $\mathsf{poly}\mkern-1.5mu\left(n\right)$ means some polynomial in $n$.
II Problem Formulation
Let $\mathcal{X}$ be a finite alphabet with $|\mathcal{X}|=d\geq 2$ and consider the set $\mathcal{P}_{\varepsilon}=\left\{P\in\mathcal{P}(\mathcal{X})\,\middle|\,\forall\,x\in\mathcal{X},\ P(x)\geq\varepsilon\right\}$ for some $\varepsilon>0$. Note that $\mathcal{P}_{\varepsilon}$ is compact and chosen to ensure that the KL divergences between these distributions are bounded and uniformly continuous. The underlying distributions are described by a pair of distinct distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}=\left\{(P,Q)\,\middle|\,P,Q\in\mathcal{P}_{\varepsilon},\ P\neq Q\right\}$, and $(P_{0},P_{1})$ is unknown to the decision maker. Under ground truth $\theta\in\{0,1\}$, the decision maker observes i.i.d. testing samples $X_{k}$’s following $P_{\theta}$, along with i.i.d. training samples $T_{0,k}$’s and $T_{1,k}$’s following $P_{0}$ and $P_{1}$ respectively to learn about the unknown underlying distributions. Note that the testing and training samples are mutually independent. Let $n\in\mathbb{N}$ be an index of the problem, which will later be related to the number of testing samples in expectation. Let $\alpha,\beta>0$ be two fixed problem parameters indicating the asymptotic ratios of the number of training samples to $n$. The objective of the decision maker is to output $\hat{\theta}\in\{0,1\}$ as an estimation of the unknown ground truth $\theta$, based on the observed samples.
Next let us specify different setups regarding the sequentiality of taking samples.
  • $\mathsf{Fully\text{-}Sequential}$: the testing and training samples are all sequentially observed. At time $k$, there are $k$ testing samples and $N_{k}=\left\lceil\alpha k\right\rceil,M_{k}=\left\lceil\beta k\right\rceil$ training samples from each distribution. A test is a pair $\Phi_{n}=(\tau_{n},\delta_{n})$ where $\tau_{n}\in\mathbb{N}$ is a Markov stopping time with respect to the filtration $\mathcal{F}_{k}=\sigma(X^{k},T_{0}^{N_{k}},T_{1}^{M_{k}})$. We may write $\tau_{n}$ as $\tau$ when it is clear from the context. The decision rule $\delta_{n}:\mathcal{X}^{\tau}\times\mathcal{X}^{N_{\tau}}\times\mathcal{X}^{M_{\tau}}\to\{0,1\}$ is a $\mathcal{F}_{\tau}$-measurable function, and the output is denoted as $\hat{\theta}$.
  • $\mathsf{Semi\text{-}Sequential\text{-}1}$: the testing samples $X_{k}$’s arrive sequentially, while the numbers of $P_{0}$/$P_{1}$-training samples are fixed to $N=\left\lceil\alpha n\right\rceil$ and $M=\left\lceil\beta n\right\rceil$. Denote the two training sequences as $T_{0}^{N}$ and $T_{1}^{M}$. A test $\Phi_{n}=(\tau_{n},\delta_{n})$ is similarly defined as in the $\mathsf{Fully\text{-}Sequential}$  setup, with time-dependent variables $N_{k},M_{k}$ replaced by the constants $N,M$.
  • $\mathsf{Semi\text{-}Sequential\text{-}2}$: the number of the testing samples is fixed to $n$ and the training samples are sequentially observed. That is, at time $k$, there are $N_{k}=\left\lceil\alpha k\right\rceil$ and $M_{k}=\left\lceil\beta k\right\rceil$ training samples from each distribution. A test $\Phi_{n}=(\tau_{n},\delta_{n})$ is similarly defined as in the $\mathsf{Fully\text{-}Sequential}$  setup, with the number of testing samples $k$ replaced by the constant $n$.
  • $\mathsf{Fixed\text{-}Length}$: all three sequences have fixed-length. Note that it can be viewed as restricting $\tau_{n}=n$ in either one of the above sequential setups.
In the following, when $X^{k}$ is observed, denote the empirical distribution as $\hat{P}^{k}$, where $\hat{P}^{k}(x)=\frac{1}{k}\sum_{i=1}^{k}\mathbbm{1}{\left\{X_{i}=x\right\}}$ for $x\in\mathcal{X}$. The empirical distributions of $T_{0}^{N}$ and $T_{1}^{M}$ are denoted as $\hat{P}_{0}^{N}$ and $\hat{P}_{1}^{M}$ (we omit $N,M$ if it is clear from the context).
The performance of tests is measured by the error probability and the number of samples used. Given $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$ and $\theta\in\{0,1\}$, the error probability is defined as $\pi_{\theta}(\Phi_{n}|P_{0},P_{1})=\mathbb{P}_{\theta}\{\hat{\theta}\neq\theta\}$, where $\mathbb{P}_{\theta}$ is the shorthand notation for the joint probability law of the testing sequence and training sequences. The average number of samples used can be described by the expected stopping time $\mathbb{E}_{\theta}[\tau_{n}|P_{0},P_{1}]$, where the expectation is taken under $\mathbb{P}_{\theta}$. Since the underlying distributions are unknown, it is natural to ask for some universal guarantees on the performance. The universality constraints are twofold.
  • First, to compare with fixed-length tests, we set a universality constraint on the expected stopping time to be at most $n$. Let $\{\Phi_{n}\}$ be a sequence of tests where $\Phi_{n}$ satisfies $\mathbb{E}_{\theta}[\tau_{n}|P_{0},P_{1}]\leq n$ for all underlying distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$ and ground truth $\theta$. The type-I and type-II error exponents of $\{\Phi_{n}\}$ given $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$ are defined as
    $e_{\theta}(P_{0},P_{1})=\liminf\limits_{n\to\infty}\frac{-\log{\pi_{\theta}(\Phi_{n}|P_{0},P_{1})}}{n},\quad\text{for }\theta=0,1.$
  • Second, we adopt the competitive Neyman-Pearson criterion proposed in [7] and set a universality constraint on the type-I error exponent. Let $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ be a pre-set distribution-dependent constraint function. We focus on tests satisfying
    $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1}),\quad\forall\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}.$
The goal is to characterize the maximum $e_{1}(P_{0},P_{1})$ that can be achieved for tests satisfying these universality constraints and to find such a test, if possible, that achieves the maximum uniformly over all possible underlying distributions.
We close this section with the following assumption on $\lambda$ that will be used in proving the achievability in the $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Semi\text{-}Sequential\text{-}2}$ setup.
Assumption 1.
The function $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ can be extended to a continuous function $\bar{\lambda}:\mathcal{P}_{\varepsilon}\times\mathcal{P}_{\varepsilon}\to[0,\infty)$.
III Main Results
To present the results, we first introduce two divergences. The Rényi Divergence of order $\frac{\alpha}{1+\alpha}$ of $P$ from $Q$ can be expressed as
$\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P\middle\|Q\right)=\min_{V\in\mathcal{P}(\mathcal{X})}\{\alpha\mathrm{D}\mkern-1.5mu\left(V\middle\|P\right)+\mathrm{D}\mkern-1.5mu\left(V\middle\|Q\right)\},$ (1)
and the minimizer is $V^{*}$, where
$V^{*}(x)=\frac{P(x)^{\frac{\alpha}{1+\alpha}}Q(x)^{\frac{1}{1+\alpha}}}{\sum_{x^{\prime}\in\mathcal{X}}P(x^{\prime})^{\frac{\alpha}{1+\alpha}}Q(x^{\prime})^{\frac{1}{1+\alpha}}}\quad\text{for }x\in\mathcal{X}.$ (2)
The $\alpha$-weighted generalized Jensen-Shannon (GJS) divergence of $Q$ from $P$ is defined as
$\mathrm{GJS}\mkern-1.5mu\left(P,Q,\alpha\right)=\alpha\mathrm{D}\mkern-1.5mu\left(P\middle\|\frac{\alpha P+Q}{\alpha+1}\right)+\mathrm{D}\mkern-1.5mu\left(Q\middle\|\frac{\alpha P+Q}{\alpha+1}\right).$
We start with the $\mathsf{Fully\text{-}Sequential}$ setup where all the samples arrive sequentially.
Theorem 1 ($\mathsf{Fully\text{-}Sequential}$).
Given $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ and let $\{\Phi_{n}\}$ be a sequence of $\mathsf{Fully\text{-}Sequential}$ tests such that for any underlying distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$,
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, $\mathbb{E}_{\theta}[\tau_{n}|P_{0},P_{1}]\leq n\ \forall\,\theta\in\{0,1\}$,
  • $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1})$.
Then for any $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, $e_{1}(P_{0},P_{1})$ is upper bounded by
$e_{1,\mathsf{seq}}^{*}(P_{0},P_{1}):=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \kappa(P_{0},P_{1})\right\},$ (3)
where
$\kappa(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}(Q_{0},Q_{1})\in\mathcal{D}_{\varepsilon}\\ g_{1}(Q_{1},Q_{0},Q_{1})<0\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)},$
$\text{and}\quad g_{1}(Q,Q_{0},Q_{1})$ $=\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\lambda(P_{0}^{\prime},P_{1}^{\prime})\Big{)}.$
Moreover, the upper bound $e_{1,\mathsf{seq}}^{*}(P_{0},P_{1})$ can be achieved simultaneously for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$.
The two terms in (3) result from the fact that if the empirical distributions are close to some true distributions, then the test should not stop too late in order to satisfy the universality constraint on the expected stopping time. Moreover, if the test stops at around time $n$, it should output decision according to the optimal $\mathsf{Fixed\text{-}Length}$ tests in order to satisfy the universality constraint on the type-I error exponent. In particular, $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$ corresponds to the region of true distributions under ground truth 00; while $\kappa(P_{0},P_{1})$ corresponds to the region of true distributions under ground truth $1$ that falls in the $\lambda$-balls centered at distributions under ground truth 00.
Next we look at the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup where the testing samples $X_{k}$’s arrive sequentially and the numbers of $P_{0}$/$P_{1}$-training samples are fixed.
Theorem 2 ($\mathsf{Semi\text{-}Sequential\text{-}1}$).
Given $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ and let $\{\Phi_{n}\}$ be a sequence of $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests such that for any underlying distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$,
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, $\mathbb{E}_{\theta}[\tau_{n}|P_{0},P_{1}]\leq n\ \forall\,\theta\in\{0,1\}$,
  • $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1})$.
Then for any $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, $e_{1}(P_{0},P_{1})$ is upper bounded by
$e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1}):=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \kappa(P_{0},P_{1}),\ \mu(P_{0},P_{1})\right\},$ (4)
where
$\mu(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\\ g(Q_{0},Q_{1})<0\end{subarray}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)},$
$\text{and}\quad g(Q_{0},Q_{1})$ $=\inf_{P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\lambda(P_{1},P_{1}^{\prime})\Big{)}.$
Moreover, if $\lambda$ satisfies Assumption 1, then $e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})$ can be achieved simultaneously for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$.
The term $\mu(P_{0},P_{1})$ comes from the limitation by the fixed-length training sequences, which marks the difference between Theorem 2 and the $\mathsf{Fully\text{-}Sequential}$ result in Theorem 1.
The next theorem summarizes the result for the $\mathsf{Semi\text{-}Sequential\text{-}2}$ setup where the testing samples have fixed length and the training samples arrive sequentially.
Theorem 3 ($\mathsf{Semi\text{-}Sequential\text{-}2}$).
Given $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ and let $\{\Phi_{n}\}$ be a sequence of $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests such that for any underlying distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$,
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, $\mathbb{E}_{\theta}[\tau_{n}|P_{0},P_{1}]\leq n\ \forall\,\theta\in\{0,1\}$,
  • $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1})$.
Then for any $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, $e_{1}(P_{0},P_{1})$ is upper bounded by
$e_{1,\mathsf{semi2}}^{*}(P_{0},P_{1}):=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \kappa(P_{0},P_{1}),\ \nu(P_{0},P_{1})\right\}.$ (5)
where
$\nu(P_{0},P_{1})=\inf_{\begin{subarray}{c}Q\in\mathcal{P}(\mathcal{X})\\ \mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right)<\lambda(P_{0},P_{1})\end{subarray}}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right).$
Moreover, if $\lambda$ satisfies Assumption 1, then $e_{1,\mathsf{semi2}}^{*}(P_{0},P_{1})$ can be achieved simultaneously for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$.
Similar as in Theorem 2, the third term comes from the limitation limitation by the fixed-length testing sequences. Notice that this term $\nu(P_{0},P_{1})$ is exactly the optimal type-II error exponent for the binary hypothesis problem, given the type-I error exponent $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1})$. This upper bound is natural because the test cannot do better than when the underlying distributions $(P_{0},P_{1})$ are known.
Theorem 1, Theorem 2 and Theorem 3 will be proved as corollaries of a more general theorem about composite hypothesis testing problem in Section IV. Finally, for comparison, we also restate the $\mathsf{Fixed\text{-}Length}$ result in [7], restricting the underlying distributions within $\mathcal{D}_{\varepsilon}$.
Theorem 4 ($\mathsf{Fixed\text{-}Length}$[7]).
Given $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ and let $\{\Phi_{n}\}$ be a sequence of $\mathsf{Fixed\text{-}Length}$ tests such that for any underlying distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, the type-I error exponent satisfies $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1})$. Then for any $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, $e_{1}(P_{0},P_{1})$ is upper bounded by
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1}):=\inf_{\begin{subarray}{c}Q,Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\\ g_{1}(Q,Q_{0},Q_{1})<0\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)},$
Moreover, the upper bound $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$ can be achieved simultaneously for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$.
Remark 1.
For Theorem 1 and Theorem 4, the results still hold even if we replace the distribution set $\mathcal{D}_{\varepsilon}$ by $\mathcal{D}=\left\{(P,Q)\,\middle|\,\mathsf{supp}(P)=\mathsf{supp}(Q)=\mathcal{X},\ P\neq Q\right\}$.
Here we take a closer look at these results and compare the optimal achievable error exponents for $\mathsf{Fully\text{-}Sequential}$, $\mathsf{Semi\text{-}Sequential\text{-}1}$, $\mathsf{Semi\text{-}Sequential\text{-}2}$ and $\mathsf{Fixed\text{-}Length}$ settings to show the benefit of sequentiality. Observe that for any $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$,
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})\leq e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})\leq e_{1,\mathsf{seq}}^{*}(P_{0},P_{1}),$ (6)
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})\leq e_{1,\mathsf{semi2}}^{*}(P_{0},P_{1})\leq e_{1,\mathsf{seq}}^{*}(P_{0},P_{1}).$ (7)
The first inequality of (6) holds because a $\mathsf{Fixed\text{-}Length}$ test is a special case of $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests. However, $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests are not a special case of $\mathsf{Fully\text{-}Sequential}$ tests as the whole training sequences are observed even when stopping before time $n$. Instead, the second inequality of (6) is obtained by directly comparing the expressions defined in (4) and (3). Following similar arguments, we get the inequalities (7) for $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests. A natural question is whether these inequalities are strict. In the following, we analytically address the question under two representative classes of $\lambda$: $\lambda\equiv\lambda_{0}$ being a constant in Section III-A, and $\lambda$ allowing efficient universal tests in Section III-B.
III-A Constant Constraint
Take $\lambda(P_{0},P_{1})\equiv\lambda_{0}$ for some $\lambda_{0}>0$. It is clear that such $\lambda$ satisfies Assumption 1. For the $\mathsf{Fixed\text{-}Length}$ case, this is exactly the setup in Ziv [5] and Gutman [6] (generalized Neyman-Pearson criterion). The following proposition shows that $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ tests achieve the same optimal error exponents under constant constraint.
Proposition 1.
If $\lambda(P_{0},P_{1})\equiv\lambda_{0}$ for some $\lambda_{0}>0$, then $\kappa(P_{0},P_{1})\leq\mu(P_{0},P_{1})$ for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, and hence
$e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})=e_{1,\mathsf{seq}}^{*}(P_{0},P_{1})=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \kappa(P_{0},P_{1})\right\}\quad\forall\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}.$
Proposition 1 shows that there is no additional gain due to sequentiality in taking training samples in the constant constraint case. The proof is in Appendix B-A. Next we provide some numerical comparisons. In resemble to the error exponent trade-off in the binary hypothesis testing problem, choose a pair of distributions $(P_{0}^{*},P_{1}^{*})\in\mathcal{D}_{\varepsilon}$ and plot the upper bounds on the type-II error exponent versus the type-I error exponent constraint $\lambda(P_{0}^{*},P_{1}^{*})$.
Figure 2: The optimal type-II error exponents under constant type-I error exponent constraint $\lambda(P_{0},P_{1})\equiv\lambda_{0}$. Fix $\mathcal{X}=\{0,1\}$, $\varepsilon=0.01$, $\alpha=2$, and choose $P_{0}^{*}=[0.6,0.4]$, $P_{1}^{*}=[0.1,0.9]$. Let $\lambda_{0}$ increase from $0.001$ to $\mathrm{GJS}\mkern-1.5mu\left(P_{0}^{*},P_{1}^{*},\alpha\right)$ to obtain a curve. Note that $\kappa(P_{0}^{*},P_{1}^{*})$ is plot under two different values of $\beta$, whereas $e_{1,\mathsf{fix}}^{*}(P_{0}^{*},P_{1}^{*})$ and $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}^{*}\middle\|P_{1}^{*}\right)$ are independent of $\beta$.
In Figure 2, there is a trade-off between the two error exponents for every setup. For $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ tests, the optimal type-II error exponent is the minimum of the red and blue curve. Comparing to $\mathsf{Fixed\text{-}Length}$ test, the benefit of sequentiality is more significant when $\beta$ is larger. This can also be observed analytically, and the results are summarized in the following proposition.
Proposition 2.
If $\lambda(P_{0},P_{1})\equiv\lambda_{0}$ for some $\lambda_{0}>0$, then $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$ and $\kappa(P_{0},P_{1})$ can be written in the same form, with coefficient $1$ replaced by $1+\beta$:
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q,Q_{0}\in\mathcal{P}_{\varepsilon}\\ \mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q,\alpha\right)\leq\lambda_{0}\end{subarray}}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right),$
$\kappa(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}\\ \mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q_{1},\alpha\right)\leq\lambda_{0}\end{subarray}}(1+\beta)\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right).$
Moreover, for a pair of distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$:
  • If $\mathrm{GJS}\mkern-1.5mu\left(P_{0},P_{1},\alpha\right)\leq\lambda_{0}$, then $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})=e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})=e_{1,\mathsf{semi2}}^{*}(P_{0},P_{1})=e_{1,\mathsf{seq}}^{*}(P_{0},P_{1})=0$.
  • If $\mathrm{GJS}\mkern-1.5mu\left(P_{0},P_{1},\alpha\right)>\lambda_{0}$, then $\kappa(P_{0},P_{1})$ is strictly increasing in $\beta$ and
    $0<e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})<e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})=e_{1,\mathsf{seq}}^{*}(P_{0},P_{1}).$
As a result, unlike in the $\mathsf{Fixed\text{-}Length}$ setup, the training sequence $T_{1}$ can improve the error exponent in $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ setups, and the gain is more significant as $\beta$ gets larger. The proof of Proposition 2 can be found in Appendix B-B. On the other hand, Figure 2 also shows that $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests may perform strictly worse than $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ tests sometimes, depending on the value of $\beta$. Recall that $\nu(P_{0},P_{1})$ does not depend on $\alpha$ nor $\beta$.
III-B Efficient Tests
Following [7, 8], we consider a special class of tests. A sequence of tests is said to be efficient [7] or universally exponentially consistent [13] if the error probabilities decay to zero exponentially for all the underlying distributions as the number of samples goes to infinity. Under our problem formulation, the formal definition of efficient tests is provided as follows.
Definition 1 (Efficient Tests).
Given $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$ and let $\{\Phi_{n}\}$ be a sequence of tests such that for any underlying distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$,
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, $\mathbb{E}_{\theta}[\tau_{n}|P_{0},P_{1}]\leq n\ \forall\,\theta\in\{0,1\}$,
  • $e_{0}(P_{0},P_{1})\geq\lambda(P_{0},P_{1})$.
We say $\{\Phi_{n}\}$ is efficient if the type-II error exponent $e_{1}(P_{0},P_{1})>0$ for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$.
In order to have tests satisfying the universality constraints and being efficient, the constraint function $\lambda$ has to satisfy certain conditions, and this will lead to further simplification of the optimal error exponents. In the following, we will introduce a necessary condition and the consequent results. Then we state a sufficient condition and show some numerical examples.
Proposition 3.
Given $\lambda:\mathcal{D}_{\varepsilon}\to(0,\infty)$, if there exists a sequence of tests satisfying the universality constraints on the expected stopping time and the type-I error exponent, furthermore being efficient; then for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, $\lambda(P_{0},P_{1})\leq\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$, and hence $\kappa(P_{0},P_{1})=\infty$. Thus
$e_{1,\mathsf{seq}}^{*}(P_{0},P_{1})$ $=\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),$
$e_{1,\mathsf{semi1}}^{*}(P_{0},P_{1})$ $=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \mu(P_{0},P_{1})\right\},$
$e_{1,\mathsf{semi2}}^{*}(P_{0},P_{1})$ $=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \nu(P_{0},P_{1})\right\}.$
Following the proof in [8], it can be shown that the error exponents of efficient tests are upper bounded by the Rényi divergence, which imposes upper bounds on $\lambda(P_{0},P_{1})$. As a result, $g_{1}(Q_{1},Q_{0},Q_{1})\geq 0$ for all $(Q_{0},Q_{1})\in\mathcal{D}_{\varepsilon}$, and hence $\kappa(P_{0},P_{1})=\infty$. Detailed proof is provided in Appendix B-C.
Next, we proceed with a sufficient condition for efficient tests.
Proposition 4 (Sufficient Condition for Efficient Tests [7]).
Take $\xi\in(0,1)$ and let $\lambda(P_{0},P_{1})=\xi\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$, then there exists efficient $\mathsf{Fixed\text{-}Length}$ tests, i.e. $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})>0$ for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$.
This proposition occurs in [7], and the proof is omitted here. Since the $\mathsf{Fixed\text{-}Length}$ tests are a special case of any types of sequential tests, it is obvious that there also exist efficient sequential tests.
Figure 3: The optimal type-II error exponents when $\lambda(P_{0},P_{1})=\xi\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$. Fix $\mathcal{X}=\{0,1\}$, $\varepsilon=0.01$, $\alpha=\beta=0.7$, and choose $P_{0}^{*}=[0.6,0.4]$, $P_{1}^{*}=[0.1,0.9]$. Let $\xi$ increase from $0.001$ to $0.999$ to obtain a curve.
Figure 3 demonstrates a special class of $\lambda$ that permits efficient tests. As shown in [7], there is a trade-off between the type-I and type-II error exponents of $\mathsf{Fixed\text{-}Length}$ tests. Note that the trade-off curve depends on the function $\lambda(\cdot,\cdot)$. In particular, given two different constraint functions $\lambda(\cdot,\cdot)$ and $\tilde{\lambda}(\cdot,\cdot)$, even if $\lambda(P_{0}^{*},P_{1}^{*})=\tilde{\lambda}(P_{0}^{*},P_{1}^{*})$, the corresponding optimal $e_{1}(P_{0}^{*},P_{1}^{*})$ may be different. On the other hand, Proposition 3 implies that $\mathsf{Fully\text{-}Sequential}$ tests completely eradicate the trade-off, consistent with the result in [8]. Given the parameters in Figure 3, there is a trade-off for $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests.
Finally, we provide a necessary and sufficient condition on $\alpha$ and $\beta$ such that $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests can achieve the same error exponents as $\mathsf{Fully\text{-}Sequential}$ tests.
Proposition 5.
Let $\lambda(P_{0},P_{1})=\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$ for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, then (8) holds if and only if $\alpha\beta\geq 1$.
$\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)\leq\mu(P_{0},P_{1})\quad\forall\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$ (8)
The proof can be found in Appendix B-D.
Proposition 5 implies that if $\alpha\beta\geq 1$, then $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests and $\mathsf{Fully\text{-}Sequential}$ tests both achieve the optimal error exponents of efficient tests $\left(\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right),\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)\right)$ for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$. On the other hand, if $\alpha\beta<1$, then for any $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests, there exists some distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$ such that the point $\left(\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right),\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)\right)$ is not achievable, and there remains a trade-off between the two error exponents. As for $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests, we do not obtain a similar result as Proposition 5, and the comparison with $\mathsf{Fully\text{-}Sequential}$ tests can be found in Section VII-C.
IV Generalization
In [7], a more general composite hypothesis testing problem was considered, and the binary classification problem in the $\mathsf{Fixed\text{-}Length}$ setup can be viewed as a special case. To provide a unified proof of the main results in Section III, we propose a more general version of the problem in [7] so that the setups with sequentially arrived samples such as $\mathsf{Semi\text{-}Sequential\text{-}1}$, $\mathsf{Semi\text{-}Sequential\text{-}2}$, and $\mathsf{Fully\text{-}Sequential}$ can be incorporated. With the results for the generalized problem, Theorem 1, Theorem 2, and Theorem 3 can immediately be obtained as corollaries. Proofs of the generalized results are given in Section V and Section VI.
Figure 4: The setup of the generalized problem.
IV-A Formulation and Results of the Generalized Problem
Let us first formalize the general setup. For convenience, in the context of the generalized problem, some notations in Section II and III are reused and overridden to avoid introducing too many new ones. The definitions of the reused notations should be clear in the respective contexts.
Let $s\in\mathbb{N}$ be the number of sequences. For each $i=1,2,\dots,s$, $\mathcal{X}_{i}$ is a finite alphabet and $\alpha_{i}>0$ is a constant. Let $\mathscr{D}:=\mathcal{P}(\mathcal{X}_{1})\times\mathcal{P}(\mathcal{X}_{2})\times\dots\times\mathcal{P}(\mathcal{X}_{s})$ and disjoint sets $\mathscr{P}_{0},\mathscr{P}_{1}\subseteq\mathrm{int}(\mathscr{D})$ be the collections of possible underlying distributions when the ground truth is 00 and $1$, respectively. Fix some integer $0\leq\ell\leq s$, indicating the number of fixed-length sequences.
Given an index $n\in\mathbb{N}$, the decision maker observes $\ell$ fixed-length sequences $X_{1}^{N_{1}},X_{2}^{N_{2}},\dots,X_{\ell}^{N_{\ell}}$, where $N_{i}=\left\lceil\alpha_{i}n\right\rceil$ is the length of the $i$-th sequence $X_{i}^{N_{i}}=(X_{i,1},X_{i,2},\dots,X_{i,N_{i}})\in\mathcal{X}_{i}^{N_{i}}$ for $i=1,2,\dots,\ell$. In addition, some samples are sequentially observed. Specifically, for $i=\ell+1,\dots,s$, at time $k$, there are $\left\lceil\alpha_{i}k\right\rceil$ samples in the $i$-th sequence. Given the unknown ground truth $\theta\in\{0,1\}$, each sample $X_{i,k}$ of the $i$-th sequence i.i.d. follow $P_{\theta,i}$, which is the $i$-th component of the underlying distribution $\bm{P}_{\theta}=(P_{\theta,1},P_{\theta,2},\dots,P_{\theta,s})\in\mathscr{P}_{\theta}$. Note that all the sequences are assumed to be mutually independent. The objective of the decision maker is to output $\hat{\theta}\in\{0,1\}$ as an estimation of the ground truth $\theta$, based on the observed samples. A test $\Phi_{n}=(\tau_{n},\delta_{n})$ consists of a Markov stopping time with respect to the filtration $\mathcal{F}_{k}=\sigma(X_{1}^{N_{1}},\dots,X_{\ell}^{N_{\ell}},X_{\ell+1}^{\left\lceil\alpha_{\ell+1}k\right\rceil},\dots,X_{s}^{\left\lceil\alpha_{s}k\right\rceil})$ and a decision rule that maps observations to $\hat{\theta}$. Given ground truth $\theta$ and underlying distributions $\bm{P}_{\theta}$, the error probability $\pi_{\theta}(\Phi_{n}|\bm{P}_{\theta})=\mathbb{P}_{\theta}\{\hat{\theta}\neq\theta\}$, where $\mathbb{P}_{\theta}$ is the shorthand notation for the joint probability law of all the sequences. The expected stopping time is $\mathbb{E}_{\theta}[\tau|\bm{P}_{\theta}]$, where the expectation is taken under $\mathbb{P}_{\theta}$.
Again, we set a constraint on the expected stopping time to be at most $n$, that is, $\mathbb{E}_{\theta}[\tau|\bm{P}_{\theta}]\leq n$ for all $\theta=0,1$ and $\bm{P}_{\theta}\in\mathcal{P}_{\theta}$. The error exponent is defined as
$e_{\theta}(\bm{P}_{\theta})=\liminf_{n\to\infty}\frac{-\log\pi_{\theta}(\Phi_{n}|\bm{P}_{\theta})}{n}.$
We also consider the competitive Neyman-Pearson criterion, where a constraint is set on the type-I error exponent. Let $\lambda:\mathscr{P}_{0}\to(0,\infty)$, and we focus on tests satisfying $e_{0}(\bm{P}_{0})\geq\lambda(\bm{P}_{0})$ for all $\bm{P}_{0}\in\mathscr{P}_{0}$.
Before we proceed, let us take a look at the special case $\ell=s$, which is essentially the fixed-length setup in [7]. The slight difference is that there is only one sequence in their problem formulation. It is not difficult to modify the arguments in [7] to show that the optimal type-II error exponent is
$e^{*}_{1,\mathsf{fix}}(\bm{P}_{1})=\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{1}(\bm{Q})<0\end{subarray}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),$ (9)
where $\bm{Q}$ denotes the tuple $(Q_{1},Q_{2},\dots,Q_{s})$, and $g_{1}(\bm{Q})=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\Big{(}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}$. Also, there exist tests satisfying the constraint on the type-I error exponent and achieve $e^{*}_{1,\mathsf{fix}}(\bm{P}_{1})$ simultaneously for all $\bm{P}_{1}\in\mathscr{P}_{1}$. Hence, in the following treatment, we shall focus on the case $\ell<s$. The results are summarized in the following theorems.
Theorem 5 (Converse).
Let $\lambda:\mathscr{P}_{0}\to(0,\infty)$, and $\{\Phi_{n}\}$ be a sequence of tests such that
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, the expected stopping time $\mathbb{E}_{\theta}[\tau_{n}|\bm{P}_{\theta}]\leq n$ for all $\theta\in\{0,1\}$ and $\bm{P}_{\theta}\in\mathscr{P}_{\theta}$,
  • $e_{0}(\bm{P}_{0})\geq\lambda(\bm{P}_{0})$ for all $\bm{P}_{0}\in\mathscr{P}_{0}$.
Then for any $\bm{P}_{1}\in\mathscr{P}_{1}$,
$e_{1}(\bm{P}_{1})\leq e_{1}^{*}(\bm{P}_{1}):=\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}\cup\Omega(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),$ (10)
where $\Gamma_{0}=\left\{\bm{Q}\in\mathscr{P}_{1}\,\middle|\,g_{1}(\bm{Q})<0\right\}$ and
$\Omega(\bm{P}_{1})$ =$=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,g^{\bm{P}_{1}}(\bm{Q})<0\right\},$ (11)
$g^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)},$ (12)
$S(\bm{P}_{1})$ =$=\left\{\bm{Q}\in\mathscr{D}\,\middle|\,Q_{i}=P_{1,i}\ \forall\,i=\ell+1,\dots,s\right\}.$
The set $\Gamma_{0}$ consists of distributions in $\mathscr{P}_{1}$ that are “too close” to some distribution $\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}$. Here “closeness” is measured by the KL divergence, and $\lambda$ is used as the threshold deciding how close is “too close”. The set $S(\bm{P}_{1})$ is a slice where the last $s-\ell$ coordinates are restricted to be the same as $\bm{P}_{1}$. This corresponds to the situation that the distributions of the last $s-\ell$ sequences are known. The set $\Omega(\bm{P}_{1})$ consists of distributions that are too close to some distribution $\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}$, restricted to the slice $S(\bm{P}_{1})$.
Theorem 6 (Achievability).
Suppose the following two conditions hold:
  1. 1.
    $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$.
  2. 2.
    $\lambda:\mathscr{P}_{0}\to(0,\infty)$ can be extended to a continuous function $\bar{\lambda}:\mathrm{cl}(\mathscr{P}_{0})\to[0,\infty)$.
Then there exists a sequence of tests $\{\Phi_{n}\}$ satisfying the universality constraints
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, the expected stopping time $\mathbb{E}_{\theta}[\tau_{n}|\bm{P}_{\theta}]\leq n$ for all $\theta\in\{0,1\}$ and $\bm{P}_{\theta}\in\mathscr{P}_{\theta}$,
  • $e_{0}(\bm{P}_{0})\geq\lambda(\bm{P}_{0})$ for all $\bm{P}_{0}\in\mathscr{P}_{0}$,
and achieves the the following type-II error exponent simultaneously for all $\bm{P}_{1}\in\mathscr{P}_{1}$:
$e_{1}(\bm{P}_{1})\geq\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}\cup\bar{\Omega}(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),$ (13)
where
$\bar{\Omega}(\bm{P}_{1})$ =$=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,\bar{g}^{\bm{P}_{1}}(\bm{Q})\leq 0\right\},$ (14)
$\bar{g}^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\bar{\lambda}(\bm{P}_{0}^{\prime})\Big{)}.$ (15)
Observe that (10) in Theorem 5 (converse) shares the same form as (13) in Theorem 6 (achievability) except that $\Omega(\bm{P}_{1})$ is replaced by $\bar{\Omega}(\bm{P}_{1})$. In particular, $<$ in (11) is replaced by $\leq$ in (14), and $g^{\bm{P}_{1}}(\bm{Q})$ in (12) is replaced by $\bar{g}^{\bm{P}_{1}}(\bm{Q})$ in (15), where the infimum is taken over $\mathscr{P}_{0}\cap S(\bm{P}_{1})$ and $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})$ respectively. Also note that in (15), $\lambda$ is replaced by its extension $\bar{\lambda}$ so that the definition is valid. The reason behind this gap is that in the proof of Theorem 5, we assume the distributions of the last $s-\ell$ sequences are known, which is possible only if infinitely many samples are observed. On the other hand, in the proof of Theorem 6, the constructed test only observes finitely many samples, so those distributions are just approximated instead of perfectly known. As a result, the closure emerges when taking limit.
Let us now investigate when the converse and achievability can coincide. In the following lemma, we first show that if the two conditions in Theorem 6 hold, $\Omega(\bm{P}_{1})$ can be replaced with another set $\Omega^{\prime}(\bm{P}_{1})$ which differs from $\bar{\Omega}(\bm{P}_{1})$ only in substituting $\bar{g}^{\bm{P}_{1}}(\bm{Q})$ with $g^{\bm{P}_{1}}(\bm{Q})$. The proof is given in Appendix C-A.
Lemma 1.
Suppose the two conditions in Theorem 6 hold, then
$\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)=\inf_{\bm{Q}\in\Omega^{\prime}(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),$
where $\Omega^{\prime}(\bm{P}_{1})=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,g^{\bm{P}_{1}}(\bm{Q})\leq 0\right\}$.
To proceed, further note that $g^{\bm{P}_{1}}(\bm{Q})$ in (12) can be rewritten as
$g^{\bm{P}_{1}}(\bm{Q})=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\bar{\lambda}(\bm{P}_{0}^{\prime})\Big{)}.$ (16)
In general, $\mathscr{P}_{0}\cap S(\bm{P}_{1})\subseteq\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))\subseteq\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})$. Hence,
$g^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\bar{\lambda}(\bm{P}_{0}^{\prime})\Big{)}$
$\geq\bar{g}^{\bm{P}_{1}}(\bm{Q})=\inf_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\bar{\lambda}(\bm{P}_{0}^{\prime})\Big{)}.$
As we will show in Lemma 4, since the KL divergence and $\bar{\lambda}$ are continuous, taking infimum over $\mathscr{P}_{0}\cap S(\bm{P}_{1})$ in (16) is the same as taking infimum over $\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$. Thus we can obtain a sufficient condition for which the converse and achievability match, that is, $\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))=\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})$. This sufficient condition will be leveraged to prove the main results for binary classification problems in Section III.
Let us summarize the above sufficient condition in the following theorem for convenience.
Theorem 7 (Characterization of the optimal type-II error exponent).
Suppose the two conditions in Theorem 6 hold and $\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))=\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})$, then the optimal type-II error exponent of all tests that satisfy the universality constraints in Theorem 5 and Theorem 6 is $e_{1}^{*}(\bm{P}_{1})$ in (10), and it can be achieved by a sequence of tests that do not depend on the actual $\bm{P}_{1}$.
Before we proceed to the specialization to the binary classification problem, let us close this sub-section by taking a look at a special case to better understand the general result. When $\ell=0$, all the samples arrive sequentially with a fixed ratio. In this case, it can be shown that the optimal type-II error exponent can be achieved and simplified, even if $\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))\neq\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})$.
Corollary 1 (Fully-sequential composite hypothesis testing).
Suppose the two conditions in Theorem 6 hold and $\ell=0$, then the optimal type-II error exponent of all tests that satisfy the universality constraints in Theorem 5 and Theorem 6 is
$e_{1}^{*}(\bm{P}_{1})=\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),$
and it can be achieved by a sequence of tests that do not depend on the actual $\bm{P}_{1}$.
Proof.
Since $\mathscr{P}_{0}$ and $\mathscr{P}_{1}$ are disjoint, we have $S(\bm{P}_{1})=\{\bm{P}_{1}\}$ and $\mathscr{P}_{0}\cap S(\bm{P}_{1})=\emptyset$. If $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\emptyset$, simply apply Theorem 7. Otherwise, $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\{\bm{P}_{1}\}$, and thus $\bm{P}_{1}\in\mathrm{cl}(\mathscr{P}_{0})$. With the help of Lemma 4, we can observe that
$0\leq e_{1}^{*}(\bm{P}_{1})\leq\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\leq\inf_{\bm{Q}\in\mathscr{P}_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)=\inf_{\bm{Q}\in\mathrm{cl}(\mathscr{P}_{0})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)=0,$
where the last equality holds because $\bm{P}_{1}\in\mathrm{cl}(\mathscr{P}_{0})$. ∎
Remark 2.
In the fixed-length setup ($\ell=s$), the result in [7] shows matching converse and achievability without the additional conditions in Theorem 6. It is worth notice that if $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$, then the optimal type-II error exponent can be rewritten as the following:
$e^{*}_{1,\mathsf{fix}}(\bm{P}_{1})=\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{1}(\bm{Q})\leq 0\end{subarray}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
This result follows from the same argument as in Lemma 1 and some observations in Section VII.
IV-B Proof of the Results for the Binary Classification Problem
IV-B1 Proof of Theorem 1 and Theorem 2
Arrange the three sequences by $T_{0},T_{1},X$. Then the binary classification problem can be viewed as a special case of the composite hypothesis testing problem by taking $s=3$, $\mathcal{X}_{i}=\mathcal{X}$ for $i=1,2,3$, $\alpha_{1}=\alpha$, $\alpha_{2}=\beta$, and $\alpha_{3}=1$. The two hypothesis sets are
$\mathscr{P}_{0}=\left\{(P_{0},P_{1},P_{0})\,\middle|\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}\right\}\quad\text{and}\quad\mathscr{P}_{1}=\left\{(P_{0},P_{1},P_{1})\,\middle|\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}\right\}.$
For the $\mathsf{Fully\text{-}Sequential}$ setup, let $\ell=0$, and for the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup, let $\ell=2$. Note that we use $\lambda$ with an abuse of notation. We could then apply Theorem 5 and Theorem 6. It can be easily verified that $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$ for all $\bm{P}_{1}\in\mathscr{P}_{1}$ in both setups, so Theorem 7 is also applicable. Specifically, notice that $\mathrm{cl}(\mathscr{P}_{0})=\left\{(P_{0},P_{1},P_{0})\,\middle|\,P_{0},P_{1}\in\mathcal{P}_{\varepsilon}\right\}$, and the only difference from $\mathscr{P}_{0}$ is that $P_{0}$ and $P_{1}$ can be the same. For the $\mathsf{Fully\text{-}Sequential}$ setup, given any $\bm{P}_{1}=(P_{0},P_{1},P_{1})\in\mathscr{P}_{1}$, we have $P_{0}\neq P_{1}$ and hence $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\emptyset$. Moreover, one can observe that $\Omega(\bm{P}_{1})=\emptyset$ for all $\bm{P}_{1}=(P_{0},P_{1},P_{1})\in\mathscr{P}_{1}$. For the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup, given any $\bm{P}_{1}=(P_{0},P_{1},P_{1})\in\mathscr{P}_{1}$, $S(\bm{P}_{1})=\left\{(Q_{1},Q_{2},P_{1})\,\middle|\,Q_{1},Q_{2}\in\mathcal{P}(\mathcal{X})\right\}$. Observe that $\mathscr{P}_{0}\cap S(\bm{P}_{1})=\left\{(P_{1},Q_{2},P_{1})\,\middle|\,Q_{2}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}\right\}$ and $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\left\{(P_{1},Q_{2},P_{1})\,\middle|\,Q_{2}\in\mathcal{P}_{\varepsilon}\right\}$. The closure of the former set is clearly the same as the latter since $Q_{2}$ can be arbitrarily close to $P_{1}$.
The bounds in Theorem 1 and Theorem 2 can be obtained via simple calculations. The three terms $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$, $\kappa(P_{0},P_{1})$ and $\mu(P_{0},P_{1})$ correspond to the divergences between the real underlying distribution $\bm{P}_{1}=(P_{0},P_{1},P_{1})$ and the three sets $\mathscr{P}_{0}$, $\Gamma_{0}$ and $\Omega(\bm{P}_{1})$, respectively. In particular, to obtain the Rényi divergence $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$,
$\inf_{\bm{Q}\in\mathscr{P}_{0}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(P_{0}^{\prime}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(P_{1}^{\prime}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(P_{0}^{\prime}\middle\|P_{1}\right)\Big{)}$
$=\inf_{P_{0}^{\prime}\in\mathcal{P}_{\varepsilon}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(P_{0}^{\prime}\middle\|P_{0}\right)+\mathrm{D}\mkern-1.5mu\left(P_{0}^{\prime}\middle\|P_{1}\right)\Big{)}=\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right).$
The last equality follows from (1), where the Rényi divergence is expressed as a minimization problem. By Hölder’s inequality, since $P_{0},P_{1}\in\mathcal{P}_{\varepsilon}$, the minimizer described in (2) is also in $\mathcal{P}_{\varepsilon}$. Thus taking the infimum over $\mathcal{P}_{\varepsilon}$ instead of $\mathcal{P}(\mathcal{X})$ in (1) still gives the Rényi divergence. To obtain $\kappa(P_{0},P_{1})$, first observe that for $\bm{Q}=(Q_{1},Q_{2},Q_{3})$,
$g_{1}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0,1}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{0,2}^{\prime}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{0,3}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}$
$=\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}^{\prime}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{0}^{\prime}\right)-\lambda(P_{0}^{\prime},P_{1}^{\prime})\Big{)}$
[…]
As a result,
$\inf_{\bm{Q}\in\Gamma_{0}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\bm{Q}\in\mathscr{P}_{1}:g_{1}(\bm{Q})<0}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\begin{subarray}{c}(Q_{0}^{\prime},Q_{1}^{\prime})\in\mathcal{D}_{\varepsilon}\\ g_{1}(Q_{1}^{\prime},Q_{0}^{\prime},Q_{1}^{\prime})<0\end{subarray}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}^{\prime}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}^{\prime}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{1}^{\prime}\middle\|P_{1}\right)\Big{)}=\kappa(P_{0},P_{1}).$
For the $\mathsf{Semi\text{-}Sequential\text{-}1}$ setup, observe that for $\bm{Q}=(Q_{1},Q_{2},Q_{3})\in S(\bm{P}_{1})$, we have $Q_{3}=P_{1}$ and
$g^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0,1}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{0,2}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}$
$=\inf_{P_{0,2}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{0,2}^{\prime}\right)-\lambda(P_{1},P_{0,2}^{\prime})\Big{)}=g(Q_{1},Q_{2}).$
Hence,
$\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\bm{Q}\in S(\bm{P}_{1}):g^{\bm{P}_{1}}(\bm{Q})<0}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\begin{subarray}{c}Q_{1},Q_{2}\in\mathcal{P}(\mathcal{X})\\ g(Q_{0},Q_{1})<0\end{subarray}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{1}\right)\Big{)}=\mu(P_{0},P_{1}).$
IV-B2 Proof of Theorem 3
Arrange the three sequences by $X,T_{0},T_{1}$, and take $s=3$, $\ell=1$, $\mathcal{X}_{i}=\mathcal{X}$ for $i=1,2,3$, $\alpha_{1}=1$, $\alpha_{2}=\alpha$, and $\alpha_{3}=\beta$. The two hypothesis sets are
$\mathscr{P}_{0}=\left\{(P_{0},P_{0},P_{1})\,\middle|\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}\right\}\quad\text{and}\quad\mathscr{P}_{1}=\left\{(P_{1},P_{0},P_{1})\,\middle|\,(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}\right\}.$
When $\bm{P}_{1}=(P_{1},P_{0},P_{1})$, we have $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))=\{(P_{0},P_{0},P_{1})\}$. The result follows immediately from Theorem 5, Theorem 6, and Theorem 7. Specifically, $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$ and $\kappa(P_{0},P_{1})$ can be obtained using the same arguments as above with a permutation of the three sequences. As for $\nu(P_{0},P_{1})$, consider $\bm{P}_{1}=(P_{1},P_{0},P_{1})$ and observe that for $\bm{Q}=(Q_{1},Q_{2},Q_{3})\in S(\bm{P}_{1})$, we have $Q_{2}=P_{0},Q_{3}=P_{1}$. Also,
$g^{\bm{P}_{1}}(\bm{Q})=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0,1}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}=\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)-\lambda(P_{0},P_{1}),$
as there is only one element $(P_{0},P_{0},P_{1})$ in $\mathscr{P}_{0}\cap S(\bm{P}_{1})$. Thus,
$\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\bm{Q}\in S(\bm{P}_{1}):g^{\bm{P}_{1}}(\bm{Q})<0}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{2}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{3}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\begin{subarray}{c}Q_{1}\in\mathcal{P}(\mathcal{X})\\ \mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}\right)<\lambda(P_{0},P_{1})\end{subarray}}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)=\nu(P_{0},P_{1}).$
V Proof of Theorem 5
For $\bm{P}_{1}\in\mathscr{P}_{1}$, notice that the bound can be decomposed into three parts, corresponding to the sets $\mathscr{P}_{0}$, $\Gamma_{0}$ and $\Omega(\bm{P}_{1})$. We start by showing the first bound $e_{1}(\bm{P}_{1})\leq\inf_{\bm{Q}\in\mathscr{P}_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$. Let us introduce two lemmas.
Lemma 2.
Let $\bm{P}_{0}\in\mathscr{P}_{0}$, $\bm{P}_{1}\in\mathscr{P}_{1}$, and $\Phi=(\tau,\delta)$ be a test. If $\mathbb{E}_{0}\left[\tau|\bm{P}_{0}\right]<\infty$, then for any $\mathcal{E}\in\mathcal{F}_{\tau}$,
$\mathrm{d}\mkern-1.5mu\left(\mathbb{P}_{0}\{\mathcal{E}\},\mathbb{P}_{1}\{\mathcal{E}\}\right)$ $\leq\sum_{i=1}^{\ell}N_{i}\mathrm{D}\mkern-1.5mu\left(P_{0,i}\middle\|P_{1,i}\right)+\sum_{i=\ell+1}^{s}\mathbb{E}_{0}\left[\left\lceil\alpha_{i}\tau\right\rceil|\bm{P}_{0}\right]\mathrm{D}\mkern-1.5mu\left(P_{0,i}\middle\|P_{1,i}\right),$
where $\mathrm{d}\mkern-1.5mu\left(p,q\right)=p\log\frac{p}{q}+(1-p)\log\frac{1-p}{1-q}$ is the binary KL divergence.
Lemma 2 is proved by the data processing inequality of KL divergences and Wald’s identity. Details can be found in Appendix C-B.
Lemma 3.
Define the function $h:(0,1)\times(0,1)\to\mathbb{R}$ as $h(p,q)=\frac{\mathrm{d}\mkern-1.5mu\left(1-q,p\right)}{-\log p}$. Then, $\lim_{(p,q)\to(0,0)}h(p,q)=1$.
The proof is easy and stated in Appendix D-A for completeness.
Given $\Phi_{n}=(\tau_{n},\delta_{n})$, apply Lemma 2 with =$\mathcal{E}=\{\hat{\theta}=0\}$. For any $\bm{Q}\in\mathscr{P}_{0}$, by the assumption in Theorem 2, $\mathbb{E}_{0}\left[\tau_{n}|\bm{Q}\right]\leq n$ and thus
$\mathrm{d}\mkern-1.5mu\left(1-\pi_{0}(\Phi_{n}|\bm{Q}),\pi_{1}(\Phi_{n}|\bm{P}_{1})\right)$ $\leq\sum_{i=1}^{\ell}N_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+\sum_{i=\ell+1}^{s}\mathbb{E}_{0}\left[\left\lceil\alpha_{i}\tau\right\rceil|\bm{Q}\right]\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$
$\leq n\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+o(1).$
For the asymptotic behavior, observe that if $\pi_{1}(\Phi_{n}|\bm{P}_{1})$ does not vanish as $n$ goes to infinity, we simply have $e_{1}(\bm{P}_{1})=0$. So we may assume that $\lim_{n\to\infty}\pi_{1}(\Phi_{n}|\bm{P}_{1})=0$. Also by the constraint on the type-I error exponent, we know that $\pi_{0}(\Phi_{n}|\bm{P}_{0})$ goes to 00 exponentially. Using Lemma 3, we have
$e_{1}(\bm{P}_{1})\leq\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
Since this is true for all $\bm{P}_{0}\in\mathscr{P}_{0}$, the upper bound is obtained.
As for the second bound $e_{1}(\bm{P}_{1})\leq\inf_{\bm{Q}\in\Gamma_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$, denote the empirical distribution as $\hat{\bm{P}}=(\hat{P_{1}},\hat{P_{2}},\dots,\hat{P_{s}})$ where $\hat{P_{i}}$ corresponds to the $i$-th sequence for $i=1,2,\dots,s$. Observe that when $\theta=1$ and the underlying distribution is $\bm{Q}\in\mathscr{P}_{1}$, $\hat{\bm{P}}$ will be close to $\bm{Q}$ with high probability. So if the empirical distribution is close to $\bm{Q}$, the test must not stop too late in order to ensure $\mathbb{E}_{1}\left[\tau_{n}|\bm{Q}\right]\leq n$. However, stopping around $n$, if $g_{1}(\bm{Q})<0$, the test must output 00 in order to satisfy the universality constraint on the type-I error exponent according to the fixed-length result [7]. This leads to upper bounds on the type-II error exponent.
We proceed by showing the technical details. If $\Gamma_{0}\neq\emptyset$, take $\bm{Q}\in\Gamma_{0}=\left\{\bm{Q}\in\mathscr{P}_{1}\,\middle|\,g_{1}(\bm{Q})<0\right\}$. By definition, there exist some $\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}$ such that $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)<\lambda(\bm{P}_{0}^{\prime})$. Given $\epsilon>0$, define
$\mathcal{B}_{\epsilon}(\bm{Q})$ =$=\Big{\{}\bm{R}\in\mathscr{D}\,\Big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(R_{i}\middle\|Q_{i}\right)<\epsilon\Big{\}},$
$\mathcal{B}_{\lambda}^{\epsilon}(\bm{P}_{0}^{\prime})$ =$=\Big{\{}\bm{R}\in\mathscr{D}\,\Big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(R_{i}\middle\|P_{0,i}^{\prime}\right)<(1-\epsilon)\lambda(\bm{P}_{0}^{\prime})\Big{\}}.$
By the choice of $\bm{P}_{0}^{\prime}$, we can take $\epsilon$ small enough such that $\mathcal{B}_{\epsilon}(\bm{Q})\subseteq\mathcal{B}_{\lambda}^{\epsilon}(\bm{P}_{0}^{\prime})$. For each test $\Phi_{n}=(\tau_{n},\delta_{n})$, by assumption $\mathbb{E}_{1}\left[\tau_{n}|\bm{Q}\right]\leq n$. Thus by Markov’s inequality,
$\mathbb{Q}\{\tau_{n}>n\}=\mathbb{Q}\{\tau_{n}\geq n+1\}\leq\frac{\mathbb{E}_{1}\left[\tau_{n}|\bm{Q}\right]}{n+1}\leq\frac{n}{n+1}.$ (17)
Here $\mathbb{Q}$ is the shorthand notation for the joint probability law of the sequences when the underlying distribution is $\bm{Q}$.
Now look at the empirical distribution $\hat{\bm{P}}$ at time $n$. Note that even if the test stops earlier, we can still consider the samples until $n$. Let $\bm{\mathcal{P}}^{n}$ denote the collection of all possible empirical distributions corresponding to the sequences observed at time $n$. It is not hard to show that $|\bm{\mathcal{P}}^{n}|$ is polynomial in $n$. By the method of types,
$\mathbb{Q}\left\{\hat{\bm{P}}\notin\mathcal{B}_{\epsilon}(\bm{Q})\right\}$ $\leq\sum_{\hat{\bm{P}}\in\bm{\mathcal{P}}^{n}\cap\mathcal{B}_{\epsilon}(\bm{Q})^{\mathrm{c}}}2^{-\sum_{i=1}^{s}\left\lceil\alpha_{i}n\right\rceil\mathrm{D}\mkern-1.5mu\left(\hat{P}_{i}\middle\|Q_{i}\right)}$
$\leq\sum_{\hat{\bm{P}}\in\bm{\mathcal{P}}^{n}\cap\mathcal{B}_{\epsilon}(\bm{Q})^{\mathrm{c}}}2^{-n\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(\hat{P}_{i}\middle\|Q_{i}\right)}$
$\leq\mathsf{poly}\mkern-1.5mu\left(n\right)2^{-n\epsilon}.$ (18)
Combine (17) and (18), then for all $n$ large enough,
$\mathbb{Q}\left\{\hat{\bm{P}}\in\mathcal{B}_{\epsilon}(\bm{Q})\text{ and }\tau_{n}\leq n\right\}\geq 1-\frac{n}{n+1}-\mathsf{poly}\mkern-1.5mu\left(n\right)2^{-n\epsilon}\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}.$
Hence for all $n$ large enough, there exist some empirical distribution $\bm{P}^{n}=(P_{1}^{n},P_{2}^{n},\dots,P_{s}^{n})\in\bm{\mathcal{P}}^{n}\cap\mathcal{B}_{\epsilon}(\bm{Q})$ such that
$\mathbb{Q}\left\{\tau_{n}\leq n\,\middle|\,\hat{\bm{P}}=\bm{P}^{n}\right\}\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}.$ (19)
Since all the sequences in a given type class have the same probability, a key observation is that this conditional probability is independent of the underlying distributions and ground truth $\theta$. So now we consider the situation where the underlying distributions are $\bm{P}_{0}^{\prime}$ and $\theta=0$. By the method of types and the fact that $\bm{P}^{n}\in\mathcal{B}_{\epsilon}(\bm{Q})\subseteq\mathcal{B}_{\lambda}^{\epsilon}(\bm{P}_{0}^{\prime})$, for $n$ large enough,
$\mathbb{P}_{0}^{\prime}\left\{\hat{\bm{P}}=\bm{P}^{n}\right\}$ $\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}2^{-\sum_{i=1}^{s}\left\lceil\alpha_{i}n\right\rceil\mathrm{D}\mkern-1.5mu\left(P_{i}^{n}\middle\|P_{0,i}^{\prime}\right)}$
$\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}2^{-n(1-\epsilon)\lambda(\bm{P}_{0}^{\prime})}.$ (20)
Using the key observation and the inequalities (19), (20), we know that for $n$ large enough
$\mathbb{P}_{0}^{\prime}\left\{\hat{\bm{P}}=\bm{P}^{n}\text{ and }\tau_{n}\leq n\right\}\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}2^{-n(1-\epsilon)\lambda(\bm{P}_{0}^{\prime})}.$
Consider the type-I error probability
$\pi_{0}(\Phi_{n}|\bm{P}_{0}^{\prime})$ $\geq\mathbb{P}_{0}^{\prime}\left\{\hat{\theta}=1,\ \hat{\bm{P}}=\bm{P}^{n}\text{ and }\tau_{n}\leq n\right\}$
$\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}2^{-n(1-\epsilon)\lambda(\bm{P}_{0}^{\prime})}\times\mathbb{P}_{0}^{\prime}\left\{\hat{\theta}=1\,\middle|\,\hat{\bm{P}}=\bm{P}^{n}\text{ and }\tau_{n}\leq n\right\}.$
Since $\{\Phi_{n}\}$ satisfies the constraint on the type-I error exponent, that is $\pi_{0}(\Phi_{n}|\bm{P}_{0}^{\prime})\mathrel{\dot{\leq}}2^{-n\lambda(\bm{P}_{0}^{\prime})}$, we must have
$\mathbb{P}_{0}^{\prime}\left\{\hat{\theta}=1\,\middle|\,\hat{\bm{P}}=\bm{P}^{n}\text{ and }\tau_{n}\leq n\right\}=o(1).$ (21)
Again, since all the sequences in a given type class have the same probability, the above conditional probability is independent of the underlying distributions and ground truth $\theta$. When the underlying distributions are $\bm{P}_{1}$ and $\theta=1$, the error probability can be lower bounded using (19), (21) and the method of types. Specifically, for $n$ large enough,
$\pi_{1}(\Phi_{n}|\bm{P}_{1})$ $\geq\mathbb{P}_{1}\left\{\hat{\theta}=0,\ \hat{\bm{P}}=\bm{P}^{n}\text{ and }\tau_{n}\leq n\right\}$
$\geq\frac{1}{\mathsf{poly}\mkern-1.5mu\left(n\right)}2^{-n\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(P_{i}^{n}\middle\|P_{1,i}\right)}.$
Hence
$e_{1}(\bm{P}_{1})\leq\liminf_{n\to\infty}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(P_{i}^{n}\middle\|P_{1,i}\right).$
Let $\bm{P}^{\epsilon}=(P_{1}^{\epsilon},P_{2}^{\epsilon},\dots,P_{s}^{\epsilon})$ be a limit point of $\{\bm{P}^{n}\}$, then by the continuity of KL divergence,
$e_{1}(\bm{P}_{1})\leq\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(P_{i}^{\epsilon}\middle\|P_{1,i}\right).$
Also by the continuity of KL divergence, $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(P_{i}^{\epsilon}\middle\|Q_{i}\right)\leq\epsilon$. As $\epsilon$ goes to 00, $\bm{P}^{\epsilon}$ converge to $\bm{Q}$, and
$e_{1}(\bm{P}_{1})$ $\leq\liminf_{\epsilon\to 0}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(P_{i}^{\epsilon}\middle\|P_{1,i}\right)=\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
Since this hold for all $\bm{Q}\in\Gamma_{0}$, we get the desired result.
Finally for the third bound $e_{1}(\bm{P}_{1})\leq\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$, observe that the set $\Omega(\bm{P}_{1})$ corresponds to the limitation imposed by the fixed-length training sequences. The bound is obtained by a reduction from a fixed-length composite hypothesis testing problem. Intuitively, if we are allowed to drop the constraint on the expected stopping time and take infinitely many samples of the $(\ell+1)$-th to $s$-th sequences, then their underlying distributions can be fully known. Specifically, when $\theta=1$ and the underlying distribution is $\bm{P}_{1}$, this means $(P_{1,\ell+1},\dots,P_{1,s})$ are known. In this situation, consider the following equivalent problem. Suppose $(P_{1,\ell+1},\dots,P_{1,s})$ is fixed and known. The decision maker observes $\ell$ independent fixed-length sequences $X_{1}^{N_{1}},\dots,X_{\ell}^{N_{\ell}}$, and the objective is to decide between the following two hypotheses:
[…]
Recall that $S(\bm{P}_{1})=\left\{\bm{Q}\in\mathscr{D}\,\middle|\,Q_{i}=P_{1,i}\ \forall\,i=\ell+1,\dots,s\right\}$ represents the collection of distributions with the $(\ell+1)$-th to $s$-th elements equal to $P_{1,\ell+1},\dots,P_{1,s}$. Observe that this is a fixed-length composite hypothesis testing problem [^1] The case $\mathscr{P}_{0}\cap S(\bm{P}_{1})=\emptyset$ is trivial, so here we focus on the case when it is nonempty.. Given a sequence of tests $\{\Phi_{n}\}$ for the original problem, we can apply it to this new problem. Specifically, generate the $(\ell+1)$-th to $s$-th sequences using the knowledge of $(P_{1,\ell+1},\dots,P_{1,s})$. Along with the observed fixed-length sequences $X_{1}^{N_{1}},\dots,X_{\ell}^{N_{\ell}}$, the test $\Phi_{n}$ will output a decision. The above method gives a randomized test, yet theoretically it can be easily derandomized without affecting the exponential rate. One can observe a clear correspondence between the error probabilities of the two problems, and hence a correspondence between the exponential rates of error probabilities. Since $\{\Phi_{n}\}$ satisfies the universality constraints on the type-I error exponent, we can apply the result for fixed-length composite hypothesis testing (9) and get the desired bound.
VI Proof of Theorem 6
In this section we prove Theorem 6 by proposing a sequence of tests satisfying the universality constraints and achieve the desired error exponents. When $\ell<s$, the decision maker has the flexibility to decide when to stop, so intuitively we would like to take more samples before making the decision. Nevertheless, the expected stopping time should not exceed $n$, meaning the probability of taking more samples should be kept small. First consider the empirical distributions at time $n-1$, written as $\hat{\bm{P}}^{n-1}$ for simplicity. Observe that with high probability, the empirical distributions are close to the true underlying distributions. Define two subsets of $\mathscr{D}$, for $\theta=0,1$,
$\Lambda_{\theta}^{n}=\bigcup_{\bm{P}_{\theta}^{\prime}\in\mathscr{P}_{\theta}}\Big{\{}\bm{Q}\in\mathscr{D}\,\Big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{\theta,i}^{\prime}\right)<\eta_{n}\Big{\}},$
where $\eta_{n}=\big{[}2\log n+\sum_{i=1}^{s}|\mathcal{X}_{i}|\log(\left\lceil\alpha_{i}n\right\rceil+1)\big{]}/(n-1)$ is a margin vanishing in $n$. By the method of types, the empirical distributions lie in these sets with high probability. Hence the universality constraint on the expected stopping time can be satisfied with the following stopping time
$\tau_{n}=\begin{cases}n-1&\quad\text{if }\hat{\bm{P}}^{n-1}\in\Lambda_{0}^{n}\cup\Lambda_{1}^{n},\\ n^{2}&\quad\text{if }\hat{\bm{P}}^{n-1}\in\left(\Lambda_{0}^{n}\cup\Lambda_{1}^{n}\right)^{\mathrm{c}}.\end{cases}$ (22)
Here, stopping at time $n^{2}$ can be viewed as approaching “having infinitely many samples”.
To specify the decision rule, in addition to the set $\Gamma_{0}=\left\{\bm{Q}\in\mathscr{P}_{1}\,\middle|\,g_{1}(\bm{Q})<0\right\}$, let $\Gamma_{1}=\left\{\bm{Q}\in\mathscr{P}_{1}\,\middle|\,g_{1}(\bm{Q})\geq 0\right\}$. Note that $\Gamma_{0}$ consists of the distributions under ground truth $1$ that are “close enough” to some possible distributions $\bm{P}_{0}\in\mathscr{P}_{0}$ under ground truth 00. For $\theta=0,1$, let
$\mathcal{B}_{n}(\Gamma_{\theta})=$ $\bigcup_{\bm{Q}\in\Gamma_{\theta}}\Big{\{}\bm{R}\in\mathscr{D}\,\Big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(R_{i}\middle\|Q_{i}\right)<\eta_{n}\Big{\}}$
be a subset of $\mathcal{P}(\mathcal{X})^{3}$ that slightly extends $\Gamma_{\theta}$. When $\tau_{n}=n-1$, the decision rule maps the observed samples to $\hat{\theta}$, where
$\hat{\theta}=\begin{cases}0\quad\text{if }\hat{\bm{P}}^{n-1}\in\Lambda_{0}^{n}\cup\mathcal{B}_{n}(\Gamma_{0}),\\ 1\quad\text{if }\hat{\bm{P}}^{n-1}\in\mathcal{B}_{n}(\Gamma_{1}).\end{cases}$ (23)
Notice that $\mathcal{B}_{n}(\Gamma_{0})\cup\mathcal{B}_{n}(\Gamma_{1})=\Lambda_{1}^{n}$.
When $\tau_{n}=n^{2}$, the decision maker observes $(X_{1}^{N_{1}},\dots,X_{\ell}^{N_{\ell}},X_{\ell+1}^{\left\lceil\alpha_{\ell+1}n^{2}\right\rceil},\dots,X_{s}^{\left\lceil\alpha_{s}n^{2}\right\rceil})$. Denote the corresponding empirical distribution as $\hat{\bm{P}}^{n^{2}}$, and use the fixed-length test in [7] with threshold function $\lambda_{n}(\bm{P}_{0})=\lambda(\bm{P}_{0})/n$. As a result, we can ensure that the type-I error probability is of the same order as when $\tau_{n}=n-1$. Specifically, the decision rules is
$\hat{\theta}=\begin{cases}0\quad\text{if }g_{n}(\hat{\bm{P}}^{n^{2}})<0,\\ 1\quad\text{otherwise},\end{cases}$
where
$g_{n}(\bm{Q})=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}.$
Now the test is clearly defined, we are ready to analyze its performance. First we show that the proposed test satisfies the universality constraint on the expected stopping time. Given $\theta\in\{0,1\}$ and $\bm{P}_{\theta}\in\mathscr{P}_{\theta}$, we have
$\mathbb{E}_{\theta}[\tau_{n}|\bm{P}_{\theta}]$ $\leq n-1+\mathbb{P}_{\theta}\left\{\hat{\bm{P}}^{n-1}\in(\Lambda_{\theta}^{n})^{\mathrm{c}}\right\}\times n^{2}$
$\leq n-1+\prod_{i=1}^{s}(\left\lceil\alpha_{i}n\right\rceil+1)^{|\mathcal{X}_{i}|}\times 2^{-(n-1)\eta_{n}}\times n^{2}=n,$
where the last inequality follows from the method of types.
To calculate the error exponents, we introduce a useful lemma and define some sets.
Lemma 4.
Let $S\subseteq\mathbb{R}^{m}$ be a compact set, $f:S\to\mathbb{R}$ be a continuous function and $A\subseteq S$. Given $\epsilon>0$ and $x\in\mathbb{R}^{m}$, let $\mathcal{B}^{\prime}_{\epsilon}(x)=\left\{y\in\mathbb{R}^{m}:\|y-x\|<\epsilon\right\}$ be the $\epsilon$-ball centered at $x$. Let $\mathcal{B}^{\prime}_{\epsilon}(A)=\cup_{x\in A}\mathcal{B}^{\prime}_{\epsilon}(x)$. Then
  1. 1.
    $\inf_{x\in A}f(x)=\inf_{x\in\mathrm{cl}(A)}f(x)$,
  2. 2.
    $\lim_{\epsilon\to 0}\inf_{x\in\mathcal{B}^{\prime}_{\epsilon}(A)\cap S}f(x)=\inf_{x\in\mathrm{cl}(A)}f(x)$
The proof is easy and the details can be found in Appendix D-B. Since $\mathscr{D}$ can be embedded into $\mathbb{R}^{\sum_{i=1}^{s}(|\mathcal{X}_{i}|-1)}$, consider for $\theta=0,1$,
$\mathcal{B}^{\prime}_{n}(\Gamma_{\theta})=\bigcup_{\bm{Q}\in\Gamma_{\theta}}\Big{\{}\bm{R}\in\mathscr{D}\,\Big{|}\,\big{\|}\bm{R}-\bm{Q}\big{\|}<c\eta_{n}\Big{\}},$
and
$\tilde{\Lambda}_{0}^{n}=\bigcup_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\Big{\{}\bm{Q}\in\mathscr{D}\,\Big{|}\,\big{\|}\bm{Q}-\bm{P}_{0}^{\prime}\big{\|}<c\eta_{n}\Big{\}}$
By Pinsker’s inequality and some basic inequalities between 1-norm and 2-norm, we know that $\mathcal{B}_{n}(\Gamma_{\theta})\subseteq\mathcal{B}^{\prime}_{n}(\Gamma_{\theta})$ and $\Lambda_{0}^{n}\subseteq\tilde{\Lambda}_{0}^{n}$ for some constant $c>0$.
For the type-I error exponent, let $\theta=0$ and $\bm{P}_{0}\in\mathscr{P}_{0}$. Based on the stopping time, the error events can be divided into two parts. When $\tau_{n}=n-1$, there is an error only if $\hat{\bm{P}}^{n-1}\in\mathcal{B}_{n}(\Gamma_{1})$. If $\Gamma_{1}\neq\emptyset$, using the method of types,
$\mathbb{P}_{0}\left\{\hat{\bm{P}}^{n-1}\in\mathcal{B}_{n}(\Gamma_{1})\right\}$ $\leq\mathbb{P}_{0}\left\{\hat{\bm{P}}^{n-1}\in\mathcal{B}^{\prime}_{n}(\Gamma_{1})\right\}\leq\prod_{i=1}^{s}(\left\lceil\alpha_{i}n\right\rceil+1)^{|\mathcal{X}_{i}|}\times 2^{-n\tilde{\lambda}_{n}(\bm{P}_{0})},$
where
$\tilde{\lambda}_{n}(P_{0},P_{1})=\inf_{\bm{R}\in\mathcal{B}^{\prime}_{n}(\Gamma_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(R_{i}\middle\|P_{0,i}\right).$
Since $\eta_{n}$ vanishes as $n$ goes to infinity, using Lemma 4 and the definition of $\Gamma_{1}$,
$\lim_{n\to\infty}\tilde{\lambda}_{n}(\bm{P}_{0})=\inf_{\bm{R}\in\Gamma_{1}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(R_{i}\middle\|P_{0,i}\right)\geq\lambda(\bm{P}_{0}).$
When $\tau_{n}=n^{2}$, by the method of types,
$\mathbb{P}_{0}\left\{\tau_{n}=n^{2},\ \hat{\theta}=1\right\}\leq\prod_{i=1}^{\ell}(\left\lceil\alpha_{i}n\right\rceil+1)^{|\mathcal{X}_{i}|}\prod_{i=\ell+1}^{s}(\left\lceil\alpha_{i}n^{2}\right\rceil+1)^{|\mathcal{X}_{i}|}\times 2^{-n^{2}\lambda_{n}(\bm{P}_{0})}\doteq 2^{-n\lambda(\bm{P}_{0})}.$
As a result, the type-I error probability has exponential rate $\lambda(\bm{P}_{0})$, which satisfies the universality constraint on the type-I error exponent.
For the type-II error exponent, let $\theta=1$ and $\bm{P}_{1}\in\mathscr{P}_{1}$. When $\tau_{n}=n-1$, the error probability is upper bounded by
$\mathbb{P}_{1}\left\{\hat{\bm{P}}^{n-1}\in\Lambda_{0}^{n}\cup\mathcal{B}_{n}(\Gamma_{0})\right\}\leq\mathbb{P}_{1}\left\{\hat{\bm{P}}^{n-1}\in\tilde{\Lambda}_{0}^{n}\cup\mathcal{B}_{n}^{\prime}(\Gamma_{0})\right\}$ (24)
Similar as above, since $\eta_{n}$ vanishes as $n$ goes to infinity, using Lemma 4 and method of types, it can be shown that the resulting error exponent is at least $\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$. On the other hand, when $\tau_{n}=n^{2}$,
$\mathbb{P}_{1}\left\{\tau_{n}=n^{2},\ \hat{\theta}=0\right\}\leq\prod_{i=1}^{\ell}(\left\lceil\alpha_{i}n\right\rceil+1)^{|\mathcal{X}_{i}|}\prod_{i=\ell+1}^{s}(\left\lceil\alpha_{i}n^{2}\right\rceil+1)^{|\mathcal{X}_{i}|}\times 2^{-n\mu_{n}(\bm{P}_{1})},$
where
$\mu_{n}(\bm{P}_{1})=\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{n}(\bm{Q})<0\end{subarray}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\Big{)}.$
Lemma 5.
Assume $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$, and $\lambda:\mathscr{P}_{0}\to(0,\infty)$ can be extended to a continuous function $\bar{\lambda}:\mathrm{cl}(\mathscr{P}_{0})\to[0,\infty)$. Then for $\bm{P}_{1}\in\mathscr{P}_{1}$, the sequence $\{\mu_{n}(\bm{P}_{1})\}$ is non-decreasing in $n$, and
$\lim_{n\to\infty}\mu_{n}(\bm{P}_{1})\geq\inf_{\bm{Q}\in\Omega^{\prime}(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
Intuitively, as $n$ grows, to minimize $\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$, $Q_{i}$ should be close to $P_{1,i}$ for $i=\ell+1,\dots,s$, otherwise $n\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$ gets too large. Also, to have $g_{n}(\bm{Q})$ less than 00, $P_{0,i}^{\prime}$ should be close to $Q_{i}$ for $i=\ell+1,\dots,s$. So the result should approach restricting $P_{0,i}^{\prime}=Q_{i}=P_{1,i}$ for $i=\ell+1,\dots,s$. The details are given in Appendix C-C.
By Lemma 5, we know the exponential type-II error rate when $\tau_{n}=n^{2}$. Combining the above results, Theorem 6 is proved.
VII Discussion
VII-A Assumptions on $\lambda$
VII-A1 Continuity
In our achievability part, Theorem 6, it is assumed that $\lambda:\mathscr{P}_{0}\to(0,\infty)$ can be extended to a continuous function $\bar{\lambda}:\mathrm{cl}(\mathscr{P}_{0})\to[0,\infty)$, and this assumption is crucial when proving the limit of $\{\mu_{n}(\bm{P}_{1})\}$ in Lemma 5. An interesting question is what will happen for general $\lambda$. In Section VI, it is shown that our proposed test can achieve type-II error exponent
$\min\left\{\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),\ \lim_{n\to\infty}\mu_{n}(\bm{P}_{1})\right\}.$
When $\lambda$ does not satisfy the continuity assumption, there might be a gap between $\lim_{n\to\infty}\mu_{n}(\bm{P}_{1})$ and the desired quantity. In the following, we would give a lower bound on $\lim_{n\to\infty}\mu_{n}(\bm{P}_{1})$. Also, though we currently cannot improve the type-II error exponent to close the gap, we could in fact show the asymptotic optimality by increasing the type-I error exponent.
For simplicity, we introduce some notations. Define the divergence ball centered at $\bm{P}_{0}\in\mathscr{P}_{0}$ with radius $r$ as
$\mathcal{B}_{r}(\bm{P}_{0})=\big{\{}\bm{Q}\in\mathscr{D}\,\big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)<r\big{\}}.$
We use $\mathcal{B}_{\lambda}(\bm{P}_{0})$ to denote $\mathcal{B}_{\lambda(\bm{P}_{0})}(\bm{P}_{0})$ and let $\Omega_{0}=\bigcup_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\mathcal{B}_{\lambda}(\bm{P}_{0}^{\prime})$. Notice that $\Gamma_{0}=\left\{\bm{Q}\in\mathscr{P}_{1}\,\middle|\,g_{1}(\bm{Q})<0\right\}=\mathscr{P}_{1}\cap\Omega_{0}$.
Proposition 6.
If $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$ and $\lambda$ is unbounded, then $\mathscr{P}_{1}\subseteq\Gamma_{0}$ and hence $e_{1}^{*}(\bm{P}_{1})=0$ for all $\bm{P}_{1}\in\mathscr{P}_{1}$.
Proof.
For any $\bm{Q}\in\mathscr{D}$ and $\bm{P}_{0}\in\mathscr{P}_{0}$, $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)\leq C$ for some $C>0$. Since we can find $\bm{P}_{0}\in\mathscr{P}_{0}$ such that $\lambda(\bm{P}_{0})$ is arbitrarily large, we have $\mathscr{D}\subseteq\mathcal{B}_{\lambda}(\bm{P}_{0})$ and thus $\mathscr{P}_{1}\subseteq\Gamma_{0}$. ∎
By Proposition 6, unbounded constraint functions are trivial. On the other hand, given any bounded $\lambda$, we can find a substituting $\tilde{\lambda}$ that can be extended to a continuous function $\bar{\lambda}:\mathrm{cl}(\mathscr{P}_{0})\to[0,\infty)$ and does not affect $\Omega_{0}$.
Proposition 7.
Assume $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$, then given any bounded $\lambda:\mathscr{P}_{0}\to(0,\infty)$, there exist a function $\tilde{\lambda}:\mathscr{P}_{0}\to(0,\infty)$ such that
  • $\tilde{\lambda}$ can be extended to a continuous function $\bar{\lambda}:\mathrm{cl}(\mathscr{P}_{0})\to[0,\infty)$,
  • $\tilde{\lambda}(\bm{P}_{0})\geq\lambda(\bm{P}_{0})$ for all $\bm{P}_{0}\in\mathscr{P}_{0}$,
  • $\bigcup_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\mathcal{B}_{\tilde{\lambda}}(\bm{P}_{0}^{\prime})=\Omega_{0}=\bigcup_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\mathcal{B}_{\lambda}(\bm{P}_{0}^{\prime})$.
The idea is to take $\tilde{\lambda}(\bm{P}_{0})=\sup\big{\{}r>0\,\big{|}\,\mathcal{B}_{r}(\bm{P}_{0})\subseteq\Omega_{0}\big{\}}$. Detailed proof can be found in Appendix C-D. For the fixed-length setup ($\ell=s$), the optimal type-II error exponents (9) only have dependency on $\lambda$ through $\Omega_{0}$. This implies that substituting $\lambda$ with $\tilde{\lambda}$ does not affect the type-II error exponents. Similarly, for the fully-sequential setup ($\ell=0$), notice that substituting $\lambda$ with $\tilde{\lambda}$ decreases the type-II error exponents, but Corollary 1 shows the optimal type-II error exponents under $\tilde{\lambda}$ only have dependency on $\tilde{\lambda}$ through $\Gamma_{0}=\mathscr{P}_{1}\cap\Omega_{0}$. As a result, the optimal type-II error exponents under $\lambda$ is sandwiched and we can obtain a slightly stronger version of Corollary 1. Specifically, we can drop the continuity assumption on $\lambda$.
Corollary 2 (Fully-sequential composite hypothesis testing - relaxed).
Suppose $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$ and $\ell=0$, then the optimal type-II error exponent of all tests that satisfy the universality constraints in Theorem 5 and Theorem 6 is
$e_{1}^{*}(\bm{P}_{1})=\inf_{\bm{Q}\in\mathscr{P}_{0}\cup\Gamma_{0}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right),$
and it can be achieved by a sequence of tests that do not depend on the actual $\bm{P}_{1}$.
In the case of binary classification, by Proposition 7, for $\mathsf{Fixed\text{-}Length}$ and $\mathsf{Fully\text{-}Sequential}$ setups, the type-I and optimal type-II error exponents under $\tilde{\lambda}$ is at least the optimal error exponents under $\lambda$. For $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests, let $\tilde{\mu}(P_{0},P_{1})$ denote the $\mu(P_{0},P_{1})$ under $\tilde{\lambda}$. It is clear that
$\tilde{\mu}(P_{0},P_{1})\leq\lim_{n\to\infty}\mu_{n}(P_{0},P_{1})\leq\mu(P_{0},P_{1}).$
In other words, we cannot close the potential gap between $\lim_{n\to\infty}\mu_{n}(P_{0},P_{1})$ and $\mu(P_{0},P_{1})$. However we can instead increase the type-I error exponent without affecting the optimal type-II error exponents of $\mathsf{Fixed\text{-}Length}$ and $\mathsf{Fully\text{-}Sequential}$ tests, and in this case show the optimality of the corresponding type-II error exponent of $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests. The same argument works for $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests too.
Note that when there is a gap, it might be the converse bound that is not tight enough. In the proof of converse, it is assumed that we have perfect knowledge about the distribution of the testing sequence. If $\lambda$ is not continuous, any slight estimation error may cause significant difference.
VII-A2 Range
Previously we assume that $\lambda$ is a strictly positive function. Now we discuss the case where $\lambda:\mathscr{D}\to[0,\infty)$ can sometimes be 00. Let $\mathscr{Q}=\left\{\bm{P}_{0}\in\mathscr{P}_{0}\,\middle|\,\lambda(\bm{P}_{0})=0\right\}$. Then simply consider the composite hypothesis testing problem with $\mathscr{P}_{0}^{\prime}=\mathscr{P}_{0}\setminus\mathscr{Q}$ and $\mathscr{P}_{1}^{\prime}=\mathscr{P}_{1}\cup\mathscr{Q}$. The idea is as follows. Since we only care about the error exponent, for any distribution $\bm{P}_{0}\in\mathscr{Q}$, we do not have any requirement on error probability $\pi_{0}(\Phi_{n}|\bm{P}_{0})$. The only constraint to be satisfied is that $\mathbb{E}_{0}[\tau_{n}|\bm{P}_{0}]\leq n$. Putting these distributions into the alternative hypothesis does exactly what we need.
VII-B Assumptions on the Underlying Distributions
Apart from the constraint function, another possible relaxation is about the assumption on the underlying distributions. In the achievability part, it is assumed that $\mathscr{P}_{0}$ is bounded away from the boundary of $\mathscr{D}$. If we allow distributions to come from $\mathrm{int}(\mathscr{D})$, it is possible that the sequence we construct in the proof converges to some distributions that do not have full support and that causes problems such as infinite KL divergences.
In Theorem 6, additional assumptions are made in order to match the result with the converse bound in Theorem 5. Namely, a sufficient condition is that $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$. If there exist some $\bm{Q}\in\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})$ that cannot be approximated by points in $\mathscr{P}_{0}\cap S(\bm{P}_{1})$, then our current proposed tests may result in a sequence converging to that point and lead to smaller type-II error exponents. It is unclear whether it is the converse or achievability bound that is too loose. As mentioned before, as long as the number of samples are finite, the underlying distributions cannot be perfectly known. Since the error exponent is an asymptotic notion, attempts to universally achieve the converse bound may further jeopardize the performance when the sample size is small. Note that it is also possible that the assumption of perfectly known distributions in the proof of converse is too strong.
VII-C Efficient $\mathsf{Semi\text{-}Sequential\text{-}2}$ Tests
In Section III, we compare the error exponent of efficient $\mathsf{Semi\text{-}Sequential\text{-}1}$ and $\mathsf{Fully\text{-}Sequential}$ tests. Specifically, Proposition 5 provides a necessary and sufficient condition on $\alpha$ and $\beta$ such that $\mathsf{Semi\text{-}Sequential\text{-}1}$ tests can achieve the same error exponents as $\mathsf{Fully\text{-}Sequential}$ tests, and this condition is universal for all the underlying distributions. In the example in Figure 3, there is a strict gap between $\mathsf{Semi\text{-}Sequential\text{-}2}$ and $\mathsf{Fully\text{-}Sequential}$ tests. Yet this is not always the case. We know that $\nu(P_{0},P_{1})$ does not depend on $\alpha,\beta$, while $\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$ and $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$ go to 00 as $\alpha,\beta\to 0$. As a result, for any $(P_{0},P_{1})$, when $\alpha$ or $\beta$ is small enough, the $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests achieves $\left(\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right),\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)\right)$, the same as $\mathsf{Fully\text{-}Sequential}$ tests.
Next we provide a sufficinent condintion on $\alpha$ and $\beta$ such that there is a strict gap between $\mathsf{Semi\text{-}Sequential\text{-}2}$ and $\mathsf{Fully\text{-}Sequential}$ tests. First notice that for a specific pair of distributions $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, $\nu(P_{0},P_{1})$ only depends on $\lambda(P_{0},P_{1})$. Moreover, recall that the pair $\big{(}\lambda(P_{0},P_{1}),\nu(P_{0},P_{1})\big{)}$ lies on the trade-off curve of error exponents for the binary hypothesis testing problem. Classical result shows that the curve can be characterized by the parametric curve
$\left\{\big{(}\mathrm{D}\mkern-1.5mu\left(P_{\rho}\middle\|P_{0}\right),\mathrm{D}\mkern-1.5mu\left(P_{\rho}\middle\|P_{1}\right)\big{)}\,\middle|\,\rho\in[0,1]\right\},$
where $P_{\rho}$ is a tilted distribution of $P_{0}$ towards $P_{1}$, described by
$P_{\rho}(x)=\frac{P_{0}(x)^{1-\rho}P_{1}(x)^{\rho}}{\sum_{x^{\prime}\in\mathcal{X}}P_{0}(x^{\prime})^{1-\rho}P_{1}(x^{\prime})^{\rho}}\quad\text{for }x\in\mathcal{X}.$
Also, note that by (2), the Rényi divergence can be expressed by
$\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$ $=\beta\mathrm{D}\mkern-1.5mu\left(P_{\frac{\beta}{1+\beta}}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(P_{\frac{\beta}{1+\beta}}\middle\|P_{0}\right),$
$\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$ $=\alpha\mathrm{D}\mkern-1.5mu\left(P_{\frac{1}{1+\alpha}}\middle\|P_{0}\right)+\mathrm{D}\mkern-1.5mu\left(P_{\frac{1}{1+\alpha}}\middle\|P_{1}\right).$
When $\alpha=\beta=1$, we have
$\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)=\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)=\mathrm{D}\mkern-1.5mu\left(P_{\frac{1}{2}}\middle\|P_{1}\right)+\mathrm{D}\mkern-1.5mu\left(P_{\frac{1}{2}}\middle\|P_{0}\right).$
Since $\left(\mathrm{D}\mkern-1.5mu\left(P_{\frac{1}{2}}\middle\|P_{0}\right),\mathrm{D}\mkern-1.5mu\left(P_{\frac{1}{2}}\middle\|P_{1}\right)\right)$ is a point on the trade-off curve, we know that in this case the optimal error exponents of $\mathsf{Semi\text{-}Sequential\text{-}2}$ tests are strictly smaller than $\mathsf{Fully\text{-}Sequential}$ tests.
In general, to provide a necessary and sufficient condition on $\alpha,\beta$ such that there is a strict gap between $\mathsf{Semi\text{-}Sequential\text{-}2}$ and $\mathsf{Fully\text{-}Sequential}$ tests, one have to solve the equations
$\begin{cases}\mathrm{D}\mkern-1.5mu\left(P_{\rho}\middle\|P_{0}\right)=\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)\\ \mathrm{D}\mkern-1.5mu\left(P_{\rho}\middle\|P_{1}\right)=\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)\end{cases}$
to obtain a relation between $\alpha$ and $\beta$. It is unclear whether the solution depends on $(P_{0},P_{1})$. So there may not exist a necessary and sufficient condition that does not depend on $(P_{0},P_{1})$ like Proposition 5.
VII-D Different Sequential Setups for Binary Classification
Previously, we focus on $\mathsf{Fixed\text{-}Length}$, $\mathsf{Semi\text{-}Sequential\text{-}1}$, $\mathsf{Semi\text{-}Sequential\text{-}2}$, and $\mathsf{Fully\text{-}Sequential}$ setups. Nevertheless, there are more possibilities, for example $X,T_{0}$ are sequentially observed and $T_{1}$ has fixed length. In this case, the optimal type-II error exponent can be shown to be the same as that in the $\mathsf{Fully\text{-}Sequential}$ setup, $e_{1,\mathsf{seq}}^{*}(P_{0},P_{1})=\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \kappa(P_{0},P_{1})\right\}$. This means the sequentiality in taking $P_{0}$-training samples is enough, and sequentially taking $P_{1}$-training samples does not further improve the error exponents.
However, when $X,T_{1}$ are sequentially observed and $T_{0}$ has fixed length, things are a bit different. Arrange the three sequences by $T_{0},T_{1},X$, then for $\bm{P}_{1}=(P_{0},P_{1},P_{1})$, we have $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\{(P_{1},P_{1},P_{1})\}$ and $\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))=\emptyset$. Hence the condition in Theorem 6 is not satisfied. As a result, the upper bound is $e_{1,\mathsf{seq}}^{*}$, while our test is only proved to achieve
$\min\left\{\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),\ \kappa(P_{0},P_{1}),\ \inf_{Q\in\mathcal{P}(\mathcal{X}):\alpha\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)\leq\bar{\lambda}(P_{1},P_{1})}\alpha\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right)\right\}.$
If $\bar{\lambda}(P_{1},P_{1})=0$, then the third term is larger than $\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$. Notice that for efficient tests, it is guaranteed that $\bar{\lambda}(P_{1},P_{1})=0$. But in general we cannot show the optimality of the achievability result.
VII-E Multiple Hypotheses Testing
Intuitively, our results can be extended to the multiple hypotheses testing problem if the number of hypotheses is fixed. The key idea is to iteratively do binary hypothesis testing. We briefly demonstrate the process in the following. Suppose now there are $m+1$ hypotheses, that is $\theta\in\{0,1,\dots,m\}$. Let disjoint sets $\mathscr{P}_{0},\mathscr{P}_{1},\dots,\mathscr{P}_{m}\subseteq\mathrm{int}(\mathscr{D})$ be the collections of possible underlying distributions under each hypothesis. The tests, error probabilities, error exponents, and universality constraint on the expected stopping time are defined similarly as in Section IV, the only difference is that $\theta$ can be other than 00 or $1$. To extend the competitive Neyman-Pearson criterion, we consider a set of error exponent constraints $\{\lambda_{0},\lambda_{1},\dots,\lambda_{m-1}\}$, where $\lambda_{j}:\mathscr{P}_{j}\to[0,\infty)$ for $j=0,1,\dots,m-1$. We focus on tests satisfying $e_{j}(\bm{P}_{j})\geq\lambda_{j}(\bm{P}_{j})$ for all $\bm{P}_{j}\in\mathscr{P}_{j}$ and $j=0,1,\dots,m-1$. However, $\lambda_{j}$ cannot be arbitrary as there is a trade-off between the error exponents. For $j=0,1,\dots,m-1$, let $\mathscr{Q}_{j}=\left\{\bm{P}_{j}\in\mathscr{P}_{j}\,\middle|\,\lambda_{j}(\bm{P}_{j})=0\right\}$, $\mathscr{P}_{0}^{(j)}=\mathscr{P}_{j}\setminus\mathscr{Q}_{j}$, and $\mathscr{P}_{1}^{(j)}=\left(\bigcup_{k=0}^{j}\mathscr{Q}_{k}\right)\cup\left(\bigcup_{k=j+1}^{m}\mathscr{P}_{k}\right)$. For simplicity, we assume each $\mathscr{P}_{0}^{(j)}$ is bounded away from the boundary of $\mathscr{D}$. Also, $\lambda_{j}|_{\mathscr{P}_{0}^{(j)}}$ can be extended to a continuous function $\bar{\lambda}_{j}:\mathrm{cl}(\mathscr{P}_{0}^{(j)})\to[0,\infty)$. Lastly, $\mathrm{cl}(\mathscr{P}_{0}^{(j)})\cap S(\bm{P})=\mathrm{cl}(\mathscr{P}_{0}^{(j)}\cap S(\bm{P}))$ for all $\bm{P}\in\bigcup_{k=j+1}^{m}\mathscr{P}_{k}$. Again, we focus on the case $\ell<s$.
For $j=0,\dots,m-1$, consider the binary composite hypothesis testing problem with hypothesis sets $\mathscr{P}_{0}^{(j)}$, $\mathscr{P}_{1}^{(j)}$ and type-I error exponent constraint $\lambda_{j}$. Let $e^{(j+1)}:\bigcup_{k=j+1}^{m}\mathscr{P}_{k}\to[0,\infty)$ denote the corresponding optimal type-II error exponent. Specifically, for $\bm{P}\in\bigcup_{k=j+1}^{m}\mathscr{P}_{k}$,
$e^{(j+1)}(\bm{P})=\inf_{\bm{Q}\in\mathscr{P}_{0}^{(j)}\cup\Gamma_{0}^{(j)}\cup\Omega^{(j)}(\bm{P})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{i}\right),$
where
$\Gamma_{0}^{(j)}$ =$=\left\{\bm{Q}\in\mathscr{P}_{1}^{(j)}\,\middle|\,g_{1}^{(j)}(\bm{Q})<0\right\},$
$g_{n}^{(j)}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}^{(j)}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda_{j}(\bm{P}_{0}^{\prime})\Big{)},$
$\Omega^{(j)}(\bm{P})$ =$=\left\{\bm{Q}\in S(\bm{P})\,\middle|\,g^{\bm{P},(j)}(\bm{Q})<0\right\},$
$g^{\bm{P},(j)}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}^{(j)}\cap S(\bm{P})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda_{j}(\bm{P}_{0}^{\prime})\Big{)}.$
Definition 2.
A set of error exponent constraints $\{\lambda_{0},\lambda_{1},\dots,\lambda_{m-1}\}$ is feasible if for each $j=1,\dots,m-1$ and any $\bm{P}_{j}\in\mathscr{P}_{j}$,
$\lambda(\bm{P}_{j})\leq\min_{k=1,\dots,j}e^{(k)}(\bm{P}_{j}).$
Conjecture 1 (Multiple Hypotheses).
Let $\{\lambda_{0},\lambda_{1},\dots,\lambda_{m-1}\}$ be feasible, and $\{\Phi_{n}\}$ be a sequence of tests such that
  • for each $\Phi_{n}=(\tau_{n},\delta_{n})$, the expected stopping time $\mathbb{E}_{\theta}[\tau_{n}|\bm{P}_{\theta}]\leq n$ for all $\theta\in\{0,1,\dots,m\}$ and $\bm{P}_{\theta}\in\mathscr{P}_{\theta}$,
  • $e_{j}(\bm{P}_{j})\geq\lambda_{j}(\bm{P}_{j})$ for all $\bm{P}_{j}\in\mathscr{P}_{j}$ and $j=0,1,\dots,m-1$.
Then for any $\bm{P}_{m}\in\mathscr{P}_{m}$,
$e_{m}(\bm{P}_{m})\leq\min_{k=1,\dots,m}e^{(k)}(\bm{P}_{m}).$
Moreover, under the assumptions in this subsection, the following test can achieve the upper bound simultaneously for all $\bm{P}_{m}\in\mathscr{P}_{m}$. For $n\in\mathbb{N}$, let $\eta_{n}=\big{[}2\log n+\sum_{i=1}^{s}|\mathcal{X}_{i}|\log(\left\lceil\alpha_{i}n\right\rceil+1)\big{]}/(n-1)$, and for $A\in\mathscr{D}$, let
$\mathcal{B}_{n}(A)=$ $\bigcup_{\bm{Q}\in A}\Big{\{}\bm{R}\in\mathscr{D}\,\Big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(R_{i}\middle\|Q_{i}\right)<\eta_{n}\Big{\}}.$
For simplicity, for $j=0,1,\dots,m-1$, let $\mathscr{B}_{j}=\mathcal{B}_{n}(\mathscr{P}_{0}^{(j)})\cup\mathcal{B}_{n}(\Gamma_{0}^{(j)})$. The stopping time is defined as
$\tau_{n}=\begin{cases}n-1&\quad\text{if }\hat{\bm{P}}^{n-1}\in\bigcup_{j=0,\dots,m}\mathcal{B}_{n}(\mathscr{P}_{j}),\\ n^{2}&\quad\text{otherwise}.\end{cases}$
When $\tau_{n}=n-1$, the decision rule maps the observed samples to $\hat{\theta}$, where
$\hat{\theta}=\begin{cases}0\quad\text{if }\hat{\bm{P}}^{n-1}\in\mathscr{B}_{0},\\ j\quad\text{if }\hat{\bm{P}}^{n-1}\in\mathscr{B}_{j}\setminus\left(\bigcup_{k<j}\mathscr{B}_{k}\right)\text{ for some }j=1,\dots,m-1,\\ m\quad\text{otherwise}.\end{cases}$
When $\tau_{n}=n^{2}$,
$\hat{\theta}=\begin{cases}0\quad\text{if }g_{n}^{(0)}(\hat{\bm{P}}^{n^{2}})<0,\\ j\quad\text{if }g_{n}^{(j)}(\hat{\bm{P}}^{n^{2}})<0\text{ and }\min_{k<j}g_{n}^{(k)}(\hat{\bm{P}}^{n^{2}})\geq 0\text{ for some }j=1,\dots,m-1,\\ m\quad\text{otherwise}.\end{cases}$
The conjecture can probably be proved by carefully applying Theorem 7 for $m$ times. Notice that problems like outlier detection or multiple classification are special cases. However, now the expressions are way too complicated. It may require some efforts to simplify the bounds to obtain more meaningful results in specific special cases, for example $\ell=0$.
VIII Conclusion
In this work, we consider a unified framework for binary classification with universality constraints on the expected stopping time and the type-I error exponent. Furthermore, we extend to a more general composite hypothesis testing problem and provide upper and lower bounds on the error exponents, along with a sufficient condition such that the bounds match. Applying the results to binary classification, we compare the optimal error exponents in different sequential setups in a fair manner to show the benefit of sequentiality.
Throughout this work, we focus on distributions on fixed and finite alphabets. It is interesting whether this kind of tests can be extended to more general parametric distributions such as exponential families. Since the method of types is heavily used in this work and is only applicable for distributions on finite support, an alternative tool for continuous distributions is required. Another more practical direction is to consider data with high dimension or growing alphabet size. In practice, our test is difficult to compute as it involves solving a lot of optimization problems. So one may attempt to design a test with lower computational complexity and simpler statistics. Also, our focus is on the error exponents, which is an asymptotic notion. Although the test is asymptotically optimal, there is no performance guarantee for small sample size. A possible direction is to study the sample complexity under the assumption that the two underlying distributions are $\varepsilon$-separated (total variation at least $\varepsilon$), as in [14].
Appendix A Useful Tools
Here we introduce some useful lemmas that will be used throughout the following proofs. These are well-known results, so the proofs are omitted.
Lemma 6 (Optimal error exponent trade-off of binary hypothesis testing).
Let $P_{0},P_{1}\in\mathrm{int}(\mathcal{P}(\mathcal{X}))$ be two distinct distributions, and $e_{0}\in(0,\mathrm{D}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)]$ be a constant. Then
$e_{1}(e_{0}):=\inf_{Q\in\mathcal{P}(\mathcal{X}):\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right)<e_{0}}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)=\inf_{Q\in\mathcal{P}(\mathcal{X}):\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right)\leq e_{0}}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right),$
and the infimum is attained by $Q=P_{\rho}$ for some unique $\rho\in[0,1]$, where $P_{\rho}$ is a tilted distribution of $P_{0}$ towards $P_{1}$, described by
$P_{\rho}(x)=\frac{P_{0}(x)^{1-\rho}P_{1}(x)^{\rho}}{\sum_{x^{\prime}\in\mathcal{X}}P_{0}(x^{\prime})^{1-\rho}P_{1}(x^{\prime})^{\rho}}\quad\text{for }x\in\mathcal{X}.$
Moreover, if $P_{0},P_{1}\in\mathcal{P}_{\varepsilon}$, then by Hölder’s inequality, $P_{\rho}\in\mathcal{P}_{\varepsilon}$, thus
$e_{1}(e_{0})=\inf_{Q\in\mathcal{P}_{\varepsilon}:\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right)<e_{0}}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)=\inf_{Q\in\mathcal{P}_{\varepsilon}:\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right)\leq e_{0}}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right),$
Lastly, consider the parametric curve $(e_{0},e_{1})=\big{(}\mathrm{D}\mkern-1.5mu\left(P_{\rho}\middle\|P_{0}\right),\mathrm{D}\mkern-1.5mu\left(P_{\rho}\middle\|P_{1}\right)\big{)}$ for $\rho\in[0,1]$. It can be shown that $\frac{\mathrm{d}e_{1}}{\mathrm{d}e_{0}}<0$ for $\rho\in(0,1)$, and $\frac{\mathrm{d}e_{1}}{\mathrm{d}e_{0}}=0$ at $\rho=1$.
Lemma 7.
Let $I$ be an index set and $A_{i}$ be a set for each $i\in I$. Let $A=\bigcup_{i\in I}A_{i}$ and $f:A\to[0,\infty)$ be a function. Then
$\inf_{x\in A}f(x)=\inf_{i\in I}\left(\inf_{x\in A_{i}}f(x)\right).$
Appendix B Proofs of Binary Classification
B-A Proof of Proposition 1
When $\lambda(P_{0},P_{1})\equiv\lambda_{0}$, we have
$g(Q_{0},Q_{1})$ $=\inf_{P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\lambda(P_{1},P_{1}^{\prime})\Big{)}\geq\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)-\lambda_{0}$
and hence
$\mu(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\\ g(Q_{0},Q_{1})<0\end{subarray}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}\geq\inf_{\begin{subarray}{c}Q_{0}\in\mathcal{P}(\mathcal{X})\\ \alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)<\lambda_{0}\end{subarray}}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)$ (25)
Moreover,
$\kappa(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}(Q_{0},Q_{1})\in\mathcal{D}_{\varepsilon}\\ g_{1}(Q_{1},Q_{0},Q_{1})<0\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}\leq\inf_{\begin{subarray}{c}Q_{0}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}\\ g_{1}(P_{1},Q_{0},P_{1})<0\end{subarray}}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right),$ (26)
and
$g_{1}(P_{1},Q_{0},P_{1})$ $=\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(P_{1}\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(P_{1}\middle\|P_{1}^{\prime}\right)\Big{)}-\lambda_{0}\leq\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)-\lambda_{0}.$
With the help of Lemma 6, the proof is complete by observing that (26) is upper bounded by (25).
B-B Proof of Proposition 2
First we rewrite $e_{1,\mathsf{fix}}^{*}$. By Lemma 4,
$g_{1}(Q,Q_{0},Q_{1})$ $=\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)\Big{)}-\lambda_{0}$
$=\inf_{P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)\Big{)}-\lambda_{0}.$
For $P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}$, let $\mathcal{B}(P_{0}^{\prime},P_{1}^{\prime})=\left\{Q,Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\,\middle|\,\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)<\lambda_{0}\right\}$. By the continuity and convexity of KL divergence, we know that
$\mathrm{cl}(\mathcal{B}(P_{0}^{\prime},P_{1}^{\prime}))=\left\{Q,Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\,\middle|\,\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)\leq\lambda_{0}\right\}.$
Notice that $\left\{Q,Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\,\middle|\,g_{1}(Q,Q_{0},Q_{1})<0\right\}=\bigcup_{P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}}\mathcal{B}(P_{0}^{\prime},P_{1}^{\prime})$. By Lemma 7 and Lemma 6,
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q,Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\\ g_{1}(Q,Q_{0},Q_{1})<0\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}$
$=\inf_{P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}}\left\{\inf_{(Q,Q_{0},Q_{1})\in\mathcal{B}(P_{0}^{\prime},P_{1}^{\prime})}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}\right\}$
$=\inf_{P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}}\left\{\inf_{(Q,Q_{0},Q_{1})\in\mathrm{cl}(\mathcal{B}(P_{0}^{\prime},P_{1}^{\prime}))\cap\mathcal{P}_{\varepsilon}^{3}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}\right\}$
$=\inf_{\begin{subarray}{c}Q,Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}\\ g_{1}(Q,Q_{0},Q_{1})\leq 0\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}.$ (27)
The last equality follows from the fact that since $\mathcal{P}_{\varepsilon}^{2}$ is compact and the KL divergence is continuous in the concerned region, $g_{1}(Q,Q_{0},Q_{1})\leq 0$ if and only if $\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)\leq\lambda_{0}$ for some $P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}$.
Notice that for $Q,Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}$,
$g_{1}(Q,Q_{0},Q_{1})$ $=\inf_{P_{0}^{\prime},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)\Big{)}-\lambda_{0}$
$=\inf_{P_{0}^{\prime}\in\mathcal{P}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)\Big{)}-\lambda_{0}$
$=\mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q,\alpha\right)-\lambda_{0}.$ (28)
The last equality follows from the fact that the GJS divergence can be written as the following minimization problem $\mathrm{GJS}\mkern-1.5mu\left(P,Q,\alpha\right)=\min_{V\in\mathcal{P}(\mathcal{X})}\{\alpha\mathrm{D}\mkern-1.5mu\left(P\middle\|V\right)+\mathrm{D}\mkern-1.5mu\left(Q\middle\|V\right)\}$, with the minimizer $\frac{\alpha P+Q}{\alpha+1}$. Combining (27) and (28),
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q,Q_{0}\in\mathcal{P}_{\varepsilon}\\ \mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q,\alpha\right)\leq\lambda_{0}\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\begin{subarray}{c}Q,Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}\\ \mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q,\alpha\right)\leq\lambda_{0}\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)\Big{)}.$ (29)
Next we rewrite $\kappa(P_{0},P_{1})$. By (28), we know that for $Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}$, $g_{1}(Q_{1},Q_{0},Q_{1})=\mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q_{1},\alpha\right)-\lambda_{0}$. Hence, for any $Q\in\mathcal{P}_{\varepsilon}$, the pair $(Q,Q)$ can be approximated by $(Q_{0},Q_{1})\in\mathcal{D}_{\varepsilon}$ with $g_{1}(Q_{1},Q_{0},Q_{1})<0$, since we can simply choose $Q_{0}=Q$ and $Q_{1}\in\mathcal{P}_{\varepsilon}\setminus\{Q\}$ arbitrarily close to $Q$. By Lemma 4, we can slightly change the set where the infimum is taken over in $\kappa(P_{0},P_{1})$.
$\kappa(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}(Q_{0},Q_{1})\in\mathcal{D}_{\varepsilon}\\ g_{1}(Q_{1},Q_{0},Q_{1})<0\end{subarray}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)}$
$=\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}\\ g_{1}(Q_{1},Q_{0},Q_{1})<0\end{subarray}}(1+\beta)\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right).$
Then following similar arguments in rewriting $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$, we get the desired result
$\kappa(P_{0},P_{1})=\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}\\ \mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q_{1},\alpha\right)\leq\lambda_{0}\end{subarray}}(1+\beta)\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right).$
It is then obvious that if $\mathrm{GJS}\mkern-1.5mu\left(P_{0},P_{1},\alpha\right)\leq\lambda_{0}$, then $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})$ and $\kappa(P_{0},P_{1})$ are 00. In the following we focus on the case $\mathrm{GJS}\mkern-1.5mu\left(P_{0},P_{1},\alpha\right)>\lambda_{0}$. Notice that the infimum in (29) is attainable since $\mathcal{P}_{\varepsilon}$ is compact and the GJS divergence is continuous. Also, it is 00 only if $Q=P_{1}$ and $Q_{0}=P_{0}$. However, $\mathrm{GJS}\mkern-1.5mu\left(P_{0},P_{1},\alpha\right)>\lambda_{0}$, so we have $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})>0$. To show that $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})<\kappa(P_{0},P_{1})$ and $\kappa(P_{0},P_{1})$ is strictly increasing in $\beta$, it suffices to show that for any $b\geq 1$, the following infimum is not attained by $Q_{1}=P_{1}$:
$\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}_{\varepsilon}\\ \mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q_{1},\alpha\right)\leq\lambda_{0}\end{subarray}}b\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right).$ (30)
Suppose the infimum is attained by $(Q_{0},P_{1})$. Let $V^{*}=\frac{\alpha Q_{0}+P_{1}}{\alpha+1}$, $c_{0}=\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|V^{*}\right)$ and $c_{1}=\mathrm{D}\mkern-1.5mu\left(P_{1}\middle\|V^{*}\right)$. Since $\lambda_{0}>0$, it can be shown that $Q_{0}\neq P_{1}$. Thus $V^{*}\neq P_{1}$ and $c_{1}>0$. Also, by the above arguments, we know that $Q_{0}\neq P_{0}$. As a result, we must have $c_{0}<\alpha\mathrm{D}\mkern-1.5mu\left(P_{0}\middle\|V^{*}\right)$. For $y\geq 0$, let
$Q_{0}^{y}$ $=\mathop{\mathrm{argmin}}_{Q\in\mathcal{P}_{\varepsilon}:\alpha\mathrm{D}\mkern-1.5mu\left(Q\middle\|V^{*}\right)\leq c_{0}+y}\alpha\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{0}\right),$
$Q_{1}^{y}$ $=\mathop{\mathrm{argmin}}_{Q\in\mathcal{P}_{\varepsilon}:\mathrm{D}\mkern-1.5mu\left(Q\middle\|V^{*}\right)\leq c_{1}-y}b\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right),$
and $f(y)=b\mathrm{D}\mkern-1.5mu\left(Q_{1}^{y}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}^{y}\middle\|P_{0}\right)$. Notice that $f(0)\leq\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)$. By Lemma 6, we have
$\frac{\mathrm{d}}{\mathrm{d}y}b\mathrm{D}\mkern-1.5mu\left(Q_{1}^{y}\middle\|P_{1}\right)\Big{|}_{y=0}=0\quad\text{and}\quad\frac{\mathrm{d}}{\mathrm{d}y}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}^{y}\middle\|P_{0}\right)\Big{|}_{y=0}<0.$
Hence $f^{\prime}(0)<0$, and there exists some small $y>0$ such that $f(y)<f(0)$. Also,
$\mathrm{GJS}\mkern-1.5mu\left(Q_{0}^{y},Q_{1}^{y},\alpha\right)\leq\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}^{y}\middle\|V^{*}\right)+\mathrm{D}\mkern-1.5mu\left(Q_{1}^{y}\middle\|V^{*}\right)\leq c_{0}+y+c_{1}-y=\mathrm{GJS}\mkern-1.5mu\left(Q_{0},Q_{1},\alpha\right)\leq\lambda_{0}.$
It is clear that $Q_{1}^{y}\neq P_{1}$ and this contradicts with the assumption that the infimum in (30) is attained by $(Q_{0},P_{1})$.
Finally, to show strict gain of sequentiality, it remains to show that $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})<\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$. Let $\bar{P}=\mathop{\mathrm{argmin}}_{V\in\mathcal{P}_{\varepsilon}}\alpha\mathrm{D}\mkern-1.5mu\left(V\middle\|P_{0}\right)+\mathrm{D}\mkern-1.5mu\left(V\middle\|P_{1}\right)$. Since $\lambda(\bar{P},P_{1})=\lambda_{0}>0$, we can pick $Q=\eta P_{1}+(1-\eta)\bar{P}$, $Q_{0}=\eta P_{0}+(1-\eta)\bar{P}$ for some $\eta>0$ small enough such that $(Q,Q_{0},P_{1})\in\mathcal{B}(\bar{P},P_{1})$ and then
$e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})\leq\mathrm{D}\mkern-1.5mu\left(Q\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(P_{1}\middle\|P_{1}\right)<\mathrm{D}\mkern-1.5mu\left(\bar{P}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(\bar{P}\middle\|P_{0}\right)=\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right).$
As a remark, $e_{1,\mathsf{fix}}^{*}(P_{0},P_{1})<\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$ holds in general and does not limit to the case of constant $\lambda$.
B-C Proof of Proposition 3
Following the proof of converse in [8], we can obtain $\lambda(P_{0},P_{1})\leq e_{0}(P_{0},P_{1})\leq\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$, where the first inequality follows from the universality constraint on the type-I error exponent. Then for $(Q_{0},Q_{1})\in\mathcal{D}_{\varepsilon}$,
$g_{1}(Q_{1},Q_{0},Q_{1})$ $=\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\lambda(P_{0}^{\prime},P_{1}^{\prime})\Big{)}$
$\geq\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\Big{(}\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{0}^{\prime}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}^{\prime}\middle\|P_{0}^{\prime}\right)\Big{)}$
$\geq\inf_{(P_{0}^{\prime},P_{1}^{\prime})\in\mathcal{D}_{\varepsilon}}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}^{\prime}\right)$ (31)
$\geq 0$
Note that (B-C) follows from the fact that Rényi divergence can be written as a minimization problem. We then have $\kappa(P_{0},P_{1})=\infty$ as it is the infimum of an empty set.
B-D Proof of Proposition 5
First we show that if $\alpha\beta\geq 1$, then $\mu(P_{0},P_{1})\geq\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$. Recall that
$\mu(P_{0},P_{1})$ $=\inf_{\begin{subarray}{c}Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})\\ g(Q_{0},Q_{1})<0\end{subarray}}\Big{(}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)\Big{)},$
$g(Q_{0},Q_{1})$ $=\inf_{P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}}\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\lambda(P_{1},P_{1}^{\prime}).$
Now $\lambda(P_{0},P_{1})=\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}\middle\|P_{0}\right)$ for all $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$. Hence it is equivalent to show that if for some $Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})$,
$\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)<\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right),$ (32)
then for all $P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}$, we have
$\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)\geq\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}^{\prime}\middle\|P_{1}\right).$
Since the Rényi divergence can be written as a minimization problem (1), we upper bound the RHS of (32) by $\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)$ and get $\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)<\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)$. For all $P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}$,
$\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)$ $\geq\alpha\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)$
$\geq\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)$
$\geq\mathrm{D}_{\frac{\beta}{1+\beta}}\mkern-1.5mu\left(P_{1}^{\prime}\middle\|P_{1}\right).$
Second, we prove that when $\alpha\beta<1$, there exists $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$ such that $\mu(P_{0},P_{1})<\mathrm{D}_{\frac{\alpha}{1+\alpha}}\mkern-1.5mu\left(P_{0}\middle\|P_{1}\right)$. Specifically, we want to find $P_{1}\in\mathcal{P}_{\varepsilon}$, $P_{0},P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}$, and $Q_{0},Q_{1}\in\mathcal{P}(\mathcal{X})$ such that
$g(Q_{0},Q_{1})$ $\leq\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{1}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}^{\prime}\right)-\beta\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}^{\prime}\right)-\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)<0,$
$\mu(P_{0},P_{1})$ $=\alpha\mathrm{D}\mkern-1.5mu\left(Q_{0}\middle\|P_{0}\right)+\beta\mathrm{D}\mkern-1.5mu\left(Q_{1}\middle\|P_{1}\right)<\alpha\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{0}\right)+\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{1}\right),$
where
$V_{\alpha}$ $=\mathop{\mathrm{argmin}}_{V\in\mathcal{P}(\mathcal{X})}\left\{\alpha\mathrm{D}\mkern-1.5mu\left(V\middle\|P_{0}\right)+\mathrm{D}\mkern-1.5mu\left(V\middle\|P_{1}\right)\right\},\text{ and }V_{\beta}=\mathop{\mathrm{argmin}}_{V\in\mathcal{P}(\mathcal{X})}\left\{\beta\mathrm{D}\mkern-1.5mu\left(V\middle\|P_{1}^{\prime}\right)+\mathrm{D}\mkern-1.5mu\left(V\middle\|P_{1}\right)\right\}.$
Take $Q_{0}=V_{\alpha}$ and $Q_{1}=V_{\beta}$, then it suffices to have $\alpha\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{1}\right)<\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)$ and $\beta\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)<\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{1}\right)$. Observe that $(P_{0},\alpha)$ and $(P_{1}^{\prime},\beta)$ play symmetric roles. WLOG assume $\alpha\leq\beta$ and get $\alpha<1$. Arbitrarily pick $(P_{0},P_{1})\in\mathcal{D}_{\varepsilon}$, and now the goal is to find $P_{1}^{\prime}\in\mathcal{P}_{\varepsilon}\setminus\{P_{1}\}$ such that $\alpha\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{1}\right)<\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)<\frac{1}{\beta}\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{1}\right)$. This is possible since $\alpha<\frac{1}{\beta}$ and $V_{\alpha}\neq P_{1}$. The closed-form expression of $V_{\beta}$ can be derived and notice that it is continuous in $P_{1}^{\prime}$. Hence $\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)$ is continuous in $P_{1}^{\prime}$. When $P_{1}^{\prime}=P_{1}$, $\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)=0$. When $P_{1}^{\prime}=P_{0}$, since $\alpha\leq\beta$, it can be shown that $\mathrm{D}\mkern-1.5mu\left(V_{\alpha}\middle\|P_{1}\right)\leq\mathrm{D}\mkern-1.5mu\left(V_{\beta}\middle\|P_{1}\right)$. By the intermediate value theorem, there exists $P_{1}^{\prime}$ between $P_{1}$ and $P_{0}$ with the desired property.
Appendix C Proofs of General Composite Hypothesis Testing
C-A Proof of Lemma 1
By Lemma 4,
$g^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}$
$=\inf_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}.$
Notice that $g^{\bm{P}_{1}}(\bm{Q})<0$ if and only if $\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)<\lambda(\bm{P}_{0}^{\prime})$ for some $\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$. Also, since $\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$ is compact and the KL divergence and $\lambda$ are continuous, $g^{\bm{P}_{1}}(\bm{Q})\leq 0$ if and only if $\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)\leq\lambda(\bm{P}_{0}^{\prime})$ for some $\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$. For simplicity, for $\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))$, let
$\mathcal{B}(\bm{P}_{0}^{\prime})$ =$=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)<\lambda(\bm{P}_{0}^{\prime})\right\},$
¯$\bar{\mathcal{B}}(\bm{P}_{0}^{\prime})$ =$=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)\leq\lambda(\bm{P}_{0}^{\prime})\right\}.$
Then $\Omega(\bm{P}_{1})=\bigcup_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))}\mathcal{B}(\bm{P}_{0}^{\prime})$ and $\Omega^{\prime}(\bm{P}_{1})=\bigcup_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))}\bar{\mathcal{B}}(\bm{P}_{0}^{\prime})$. Notice that if $\lambda(\bm{P}_{0}^{\prime})>0$, then
$\inf_{\bm{Q}\in\mathcal{B}(\bm{P}_{0}^{\prime})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)=\inf_{\bm{Q}\in\bar{\mathcal{B}}(\bm{P}_{0}^{\prime})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
On the other hand, if $\lambda(\bm{P}_{0}^{\prime})=0$, then $\mathcal{B}(\bm{P}_{0}^{\prime})=\emptyset$ and ¯$\bar{\mathcal{B}}(\bm{P}_{0}^{\prime})=\{\bm{P}_{0}^{\prime}\}$. For all $\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})$, we have $\lambda(\bm{P}_{0}^{\prime})>0$ and $\bm{P}_{0}^{\prime}\in\mathcal{B}(\bm{P}_{0}^{\prime})$. Hence
$\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\leq\inf_{\bm{Q}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)=\inf_{\bm{Q}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$ (33)
By Lemma 7,
$\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))}\left(\inf_{\bm{Q}\in\mathcal{B}(\bm{P}_{0}^{\prime})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\right)$
$=\inf_{\begin{subarray}{c}\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))\\ \lambda(\bm{P}_{0}^{\prime})>0\end{subarray}}\left(\inf_{\bm{Q}\in\mathcal{B}(\bm{P}_{0}^{\prime})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\right).$
Also,
$\inf_{\bm{Q}\in\Omega^{\prime}(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)$
$=\inf_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))}\left(\inf_{\bm{Q}\in\bar{\mathcal{B}}(\bm{P}_{0}^{\prime})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\right)$
$=\min\left\{\inf_{\begin{subarray}{c}\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))\\ \lambda(\bm{P}_{0}^{\prime})>0\end{subarray}}\left(\inf_{\bm{Q}\in\bar{\mathcal{B}}(\bm{P}_{0}^{\prime})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\right),\inf_{\begin{subarray}{c}\bm{Q}\in\mathrm{cl}(\mathscr{P}_{0}\cap S(\bm{P}_{1}))\\ \lambda(\bm{Q})=0\end{subarray}}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\right\}$ (34)
$=\inf_{\bm{Q}\in\Omega(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
The last equality holds since the second term in (34) is lower bounded by (33).
C-B Proof of Lemma 2
Using data processing inequality of divergence:
$\mathrm{d}\mkern-1.5mu\left(\mathbb{P}_{0}\{\mathcal{E}\},\mathbb{P}_{1}\{\mathcal{E}\}\right)$ $\leq\mathrm{D}\mkern-1.5mu\left(\mathbb{P}_{0}\middle\|\mathbb{P}_{1}\right)|_{\mathcal{F}_{\tau}}$
$=\mathbb{E}_{0}\left[\sum_{i=1}^{\ell}\sum_{k=1}^{N_{i}}\log\frac{P_{0,i}(X_{i,k})}{P_{1,i}(X_{i,k})}+\sum_{i=\ell+1}^{s}\sum_{k=1}^{\left\lceil\alpha_{i}\tau\right\rceil}\log\frac{P_{0,i}(X_{i,k})}{P_{1,i}(X_{i,k})}\,\middle|\,\bm{P}_{0}\right]$
$=\sum_{i=1}^{\ell}N_{i}\mathrm{D}\mkern-1.5mu\left(P_{0,i}\middle\|P_{1,i}\right)+\sum_{i=\ell+1}^{s}\mathbb{E}_{0}\left[\left\lceil\alpha_{i}\tau\right\rceil|\bm{P}_{0}\right]\mathrm{D}\mkern-1.5mu\left(P_{0,i}\middle\|P_{1,i}\right),$
where the last equality is by Wald’s identity. One can easily check that the assumptions for Wald’s identity are satisfied since $\tau$ is a stopping time and $X_{i,k}$’s are i.i.d. sampled from a distribution with full support on a finite alphabet.
C-C Proof of Lemma 5
Recall that
$\mu_{n}(\bm{P}_{1})$ $=\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{n}(\bm{Q})<0\end{subarray}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\Big{)},$
$g_{n}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)},$
$S(\bm{P}_{1})$ =$=\left\{\bm{Q}\in\mathscr{D}\,\middle|\,Q_{i}=P_{1,i}\ \forall\,i=\ell+1,\dots,s\right\}$
$\bar{\Omega}(\bm{P}_{1})$ =$=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,\bar{g}^{\bm{P}_{1}}(\bm{Q})\leq 0\right\},$
$\bar{g}^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\bar{\lambda}(\bm{P}_{0}^{\prime})\Big{)},$
$\Omega(\bm{P}_{1})$ =$=\left\{\bm{Q}\in S(\bm{P}_{1})\,\middle|\,g^{\bm{P}_{1}}(\bm{Q})<0\right\},$
$g^{\bm{P}_{1}}(\bm{Q})$ $=\inf_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}\cap S(\bm{P}_{1})}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{\prime}\right)-\lambda(\bm{P}_{0}^{\prime})\Big{)}.$
For simplicity, let
$\mu(\bm{P}_{1})=\inf_{\bm{Q}\in\bar{\Omega}(\bm{P}_{1})}\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)=\inf_{\bm{Q}\in\bar{\Omega}(\bm{P}_{1})}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right).$
The second equality holds because $\bar{\Omega}(\bm{P}_{1})\subseteq S(\bm{P}_{1})$. To show that $\{\mu_{n}(\bm{P}_{1})\}$ is non-decreasing in $n$, observe that for any $\bm{Q}\in\mathscr{D}$, $\{g_{n}(\bm{Q})\}$ is non-decreasing in $n$. So
$\mu_{n}(\bm{P}_{1})$ $=\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{n}(\bm{Q})<0\end{subarray}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\Big{)}$
$\leq\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{n+1}(\bm{Q})<0\end{subarray}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\Big{)}$
$\leq\inf_{\begin{subarray}{c}\bm{Q}\in\mathscr{D}\\ g_{n+1}(\bm{Q})<0\end{subarray}}\Big{(}\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)+(n+1)\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{1,i}\right)\Big{)}=\mu_{n+1}(\bm{P}_{1}).$
If $\{\mu_{n}(\bm{P}_{1})\}$ is unbounded, the statement is trivially true. Otherwise, by the monotone convergence theorem, $\{\mu_{n}(\bm{P}_{1})\}$ converges. Denote the limit by $\lim_{n\to\infty}\mu_{n}(\bm{P}_{1})=E$. For any $n$, there exist $\bm{Q}^{n}\in\mathscr{D}$ and $\bm{P}_{0}^{n}\in\mathscr{P}_{0}$ such that
$\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}^{n}\middle\|P_{1,i}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}^{n}\middle\|P_{1,i}\right)\leq\mu_{n}(\bm{P}_{1})+\frac{1}{n}\leq E+\frac{1}{n},$ (35)
$\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}^{n}\middle\|P_{0,i}^{n}\right)+n\sum_{i=\ell+1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}^{n}\middle\|P_{0,i}^{n}\right)-\lambda(\bm{P}_{0}^{n})<0.$ (36)
Let $\bar{\bm{Q}}\in\mathscr{D}$ and $\bar{\bm{P}}_{0}\in\mathrm{cl}(\mathscr{P}_{0})$ be limit points of the sequences $\{\bm{Q}^{n}\}$ and $\{\bm{P}_{0}^{n}\}$. For $i=\ell+1,\dots,s$, by (35), $\mathrm{D}\mkern-1.5mu\left(Q_{i}^{n}\middle\|P_{1,i}\right)\leq\frac{1}{n}(E+\frac{1}{n})\to 0$ as $n\to\infty$, and hence $\bar{Q}_{i}=P_{1,i}$. Since $\lambda$ can be extended to a continuous function $\bar{\lambda}:\mathrm{cl}(\mathscr{P}_{0})\to[0,\infty)$ and $\mathrm{cl}(\mathscr{P}_{0})$ is compact, we know that $\bar{\lambda}$ is bounded, and so is $\lambda$. By (36), it can be shown similarly that $\bar{Q}_{i}=\bar{P}_{0,i}$, thus $\bar{P}_{0,i}=P_{1,i}$. As a result, we have $\bar{\bm{Q}},\bar{\bm{P}}_{0}\in S(\bm{P}_{1})$. If $\mathrm{cl}(\mathscr{P}_{0})\cap S(\bm{P}_{1})=\emptyset$, then $\bar{\bm{P}}_{0}$ makes a contradiction. Otherwise, by the continuity of KL divergence and $\bar{\lambda}$, we have
$\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(\bar{Q}_{i}\middle\|P_{1,i}\right)\leq E,$
$\bar{g}^{\bm{P}_{1}}(\bar{\bm{Q}})\leq$ $\sum_{i=1}^{\ell}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(\bar{Q}_{i}\middle\|\bar{P}_{0,i}\right)-\bar{\lambda}(\bar{\bm{P}}_{0})\leq 0.$
If $\bar{\Omega}(\bm{P}_{1})=\emptyset$, there is again a direct contradiction. Otherwise $E\geq\mu(\bm{P}_{0})$ by the definition of $\mu(\bm{P}_{0})$.
C-D Proof of Proposition 7
Recall that the divergence ball centered at $\bm{P}_{0}\in\mathscr{P}_{0}$ with radius $r$ is
$\mathcal{B}_{r}(\bm{P}_{0})=\big{\{}\bm{Q}\in\mathscr{D}\,\big{|}\,\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)<r\big{\}}.$
Also, $\Omega_{0}=\bigcup_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\mathcal{B}_{\lambda}(\bm{P}_{0}^{\prime})$, where $\mathcal{B}_{\lambda}(\bm{P}_{0})$ denotes $\mathcal{B}_{\lambda(\bm{P}_{0})}(\bm{P}_{0})$ for simplicity.
If $\Omega_{0}=\mathscr{D}$, it is done by simply taking $\tilde{\lambda}(\bm{P}_{0})=\sup_{\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}}\lambda(\bm{P}_{0}^{\prime})$ as $\lambda$ is bounded. If $\Omega_{0}\neq\mathscr{D}$, let $\bar{\lambda}(\bm{P}_{0})=\sup\big{\{}r\geq 0\,\big{|}\,\mathcal{B}_{r}(\bm{P}_{0})\subseteq\Omega_{0}\big{\}}$ for $\bm{P}_{0}\in\mathrm{cl}(\mathscr{P}_{0})$. The function $\bar{\lambda}$ is well-defined because there exists $\bm{Q}\in\mathscr{D}\setminus\Omega_{0}$ and $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)$ is bounded as $\mathrm{cl}(\mathscr{P}_{0})\subseteq\mathrm{int}(\mathscr{D})$. Take $\tilde{\lambda}=\bar{\lambda}|_{\mathscr{P}_{0}}$, and we first show that $\mathcal{B}_{\tilde{\lambda}}(\bm{P}_{0})\subseteq\Omega_{0}$ for each $\bm{P}_{0}\in\mathscr{P}_{0}$. Suppose there exists $\bm{Q}\in\mathcal{B}_{\tilde{\lambda}}(\bm{P}_{0})\setminus\Omega_{0}$, then we have $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)<\tilde{\lambda}(\bm{P}_{0})$. By definition, there exists $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)<r\leq\tilde{\lambda}(\bm{P}_{0})$ such that $\mathcal{B}_{r}(\bm{P}_{0})\subseteq\Omega_{0}$. Hence $\bm{Q}\in\mathcal{B}_{r}(\bm{P}_{0})\subseteq\Omega_{0}$, which makes a contradiction. It remains to show that $\bar{\lambda}$ is continuous. For any $\bm{P}_{0}\in\mathrm{cl}(\mathscr{P}_{0})$, suppose $\bar{\lambda}$ is not continuous at $\bm{P}_{0}$. Then there exists $\epsilon>0$ such that $\forall\,m\in\mathbb{N}$, $\exists\,\bm{P}_{0}^{m}\in\mathrm{cl}(\mathscr{P}_{0})$ with $\left\lVert\bm{P}_{0}^{m}-\bm{P}_{0}\right\rVert<\frac{1}{m}$ and $|\bar{\lambda}(\bm{P}_{0}^{m})-\bar{\lambda}(\bm{P}_{0})|\geq\epsilon$.
Case 1.
For infinitely many $m$, $\bar{\lambda}(\bm{P}_{0}^{m})\leq\bar{\lambda}(\bm{P}_{0})-\epsilon$. Let $r=\bar{\lambda}(\bm{P}_{0})-\epsilon/2$ and focus on the infinite subsequence. For each $m$ in the subsequence, there exists $\bm{Q}^{m}\in\mathcal{B}_{r}(\bm{P}_{0}^{m})\cap\Omega_{0}^{\mathrm{c}}$. Let $\bar{\bm{Q}}$ be a limit point of $\{\bm{Q}^{m}\}$. Since $\mathcal{B}_{\lambda}(\bm{P}_{0}^{\prime})$ is open for all $\bm{P}_{0}^{\prime}\in\mathscr{P}_{0}$, we know that $\Omega_{0}$ is open and hence $\Omega_{0}^{\mathrm{c}}$ is closed (and also bounded). Therefore $\bar{\bm{Q}}\in\Omega_{0}^{\mathrm{c}}$. However, we have $\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}^{m}\middle\|P_{0,i}^{m}\right)<r$ for each $m$ in the subsequence and by continuity of KL divergence,
$\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(\bar{Q}_{i}\middle\|P_{0,i}\right)\leq r<\bar{\lambda}(\bm{P}_{0})-\epsilon/4.$
This implies $\bar{\bm{Q}}\in\mathcal{B}_{\bar{\lambda}(\bm{P}_{0})-\epsilon/4}(\bm{P}_{0})\subseteq\Omega_{0}$, which leads to contradiction.
Case 2.
For infinitely many $m$, $\bar{\lambda}(\bm{P}_{0}^{m})\geq\bar{\lambda}(\bm{P}_{0})+\epsilon$. Let $r=\bar{\lambda}(\bm{P}_{0})+\epsilon/2$ and focus on the infinite subsequence. For each $m$ in the subsequence, $\mathcal{B}_{r}(\bm{P}_{0}^{m})\subseteq\Omega_{0}$. Take $r^{\prime}=\bar{\lambda}(\bm{P}_{0})+\epsilon/4$ and notice that for each $i=1,\dots,s$,
$\big{|}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{m}\right)-\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)\big{|}\leq\sum_{x\in\mathcal{X}_{i}}\big{|}\log P_{0,i}^{m}(x)-\log P_{0,i}(x)\big{|}.$
Since $\bm{P}_{0}^{m}\to\bm{P}_{0}$, we can find $m$ large enough such that for all $\bm{Q}\in\mathcal{B}_{r^{\prime}}(\bm{P}_{0})$,
$\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}^{m}\right)\leq\sum_{i=1}^{s}\alpha_{i}\mathrm{D}\mkern-1.5mu\left(Q_{i}\middle\|P_{0,i}\right)+\epsilon/8<r.$
Thus $\mathcal{B}_{r^{\prime}}(\bm{P}_{0})\subseteq\mathcal{B}_{r}(\bm{P}_{0}^{m})\subseteq\Omega$, leading to contradiction.
Appendix D Proofs of Technical Tools
D-A Proof of Lemma 3
Given any $\epsilon>0$, we aim to find $\delta>0$ such that for all $p,q\in(0,\delta)$, $|h(p,q)-1|\leq\epsilon$. For $p,q\in(0,\delta)$, if $\delta<0.5$,
$\mathrm{d}\mkern-1.5mu\left(1-p,q\right)$ $\geq\mathrm{d}\mkern-1.5mu\left(1-\delta,q\right)$
$=(1-\delta)\log\frac{1-\delta}{q}+\delta\log\frac{\delta}{1-q}$
$=(1-\delta)\log\frac{1}{q}+(1-\delta)\log(1-\delta)+\delta\log\delta+\delta\log\frac{1}{1-q}$
$\geq(1-\delta)\log\frac{1}{q}-\frac{2}{e\ln 2}\qquad\text{(since }\forall\,x\in(0,1),\;x\log x\geq-\frac{2}{e\ln 2}\text{)}$
For the upper bound,
$\mathrm{d}\mkern-1.5mu\left(1-p,q\right)=(1-p)\log\frac{1-p}{q}+p\log\frac{p}{1-q}\leq\log\frac{1}{q}+\delta\log\frac{1}{1-\delta}.$
Hence
$\frac{\mathrm{d}\mkern-1.5mu\left(1-p,q\right)}{\log\frac{1}{q}}\geq 1-\delta-\frac{2}{e\ln 2}\frac{1}{\log\frac{1}{q}}\geq 1-\delta-\frac{2}{e\ln 2}\frac{1}{\log\frac{1}{\delta}},$
and
$\frac{\mathrm{d}\mkern-1.5mu\left(1-p,q\right)}{\log\frac{1}{q}}\leq 1+\frac{\delta\log\frac{1}{1-\delta}}{\log\frac{1}{q}}\leq 1+\frac{\delta\log\frac{1}{1-\delta}}{\log\frac{1}{\delta}}.$
It suffices to choose $\delta>0$ small enough such that
$\delta-\frac{2}{e\ln 2}\frac{1}{\log\delta}<\epsilon\text{ and }\frac{\delta\log(1-\delta)}{\log\delta}<\epsilon.$
D-B Proof of Lemma 4
For the first part, clearly, $\inf_{x\in A}f(x)\geq\inf_{x\in\mathrm{cl}(A)}f(x)$. For $y\in\mathrm{cl}(A)$, choose a sequence $\{y_{n}\}$ such that $y_{n}\in A$ and $y_{n}\to y$. By the continuity of $f$ and the definition of infimum, we have
$f(y)=\lim_{n\to\infty}f(y_{n})\geq\inf_{x\in A}f(x).$
Since this holds for all $y\in\mathrm{cl}(A)$, it follows that $\inf_{x\in\mathrm{cl}(A)}f(x)\geq\inf_{x\in A}f(x)$.
For the second part, observe that for any $\epsilon>0$, we have $\inf_{x\in\mathcal{B}^{\prime}_{\epsilon}(A)\cap S}f(x)\leq\inf_{x\in\mathrm{cl}(A)}f(x)<\infty$. Also, $\inf_{x\in\mathcal{B}^{\prime}_{\epsilon}(A)\cap S}f(x)$ is non-decreasing as $\epsilon\to 0$. Hence we know the limit exists. Assume
$\lim_{\epsilon\to 0}\inf_{x\in\mathcal{B}^{\prime}_{\epsilon}(A)\cap S}f(x)=a<\inf_{x\in\mathrm{cl}(A)}f(x).$
For each $n\in\mathbb{N}$, there exists $y_{n}\in\mathcal{B}^{\prime}_{1/n}(A)\cap S$ such that $f(y_{n})\leq\inf_{x\in\mathcal{B}^{\prime}_{1/n}(A)\cap S}f(x)+1/n$. Since $y_{n}\in\mathcal{B}^{\prime}_{1/n}(A)$, there exists $x_{n}\in A$ such that $\left\lVert y_{n}-x_{n}\right\rVert\leq 1/n$. Consider a convergent subsequence $\{y_{n_{i}}\}$ and let $y\in S$ denote the limit point. Since $\left\lVert y-x_{n_{i}}\right\rVert\leq\left\lVert y-y_{n_{i}}\right\rVert+\left\lVert y_{n_{i}}-x_{n_{i}}\right\rVert$ converges to 00 as $i$ goes to infinity, it follows that $y\in\mathrm{cl}(A)$. By the continuity of $f$, we have
$f(y)=\lim_{i\to\infty}f(y_{n_{i}})\leq a<\inf_{x\in\mathrm{cl}(A)}f(x),$
which makes a contradiction.